### v1

# Weighted Envy Freeness With Limited Subsidies

Abstract
We explore solutions for fairly allocating indivisible items among agents assigned weights representing their entitlements. Our fairness goal is weighted-envy-freeness (WEF), where each agent deems their allocated portion relative to their entitlement at least as favorable as any other’s relative to their own. In many cases, achieving WEF necessitates monetary transfers, which can be modeled as third-party subsidies. The goal is to attain WEF with bounded subsidies.
Previous work in the unweighted setting of subsidies relied on basic characterizations of EF that fail in the weighted settings. This makes our new setting challenging and theoretically intriguing. We present polynomial-time algorithms that compute WEF-able allocations with an upper bound on the subsidy per agent in three distinct additive valuation scenarios: (1) general, (2) identical, and (3) binary. When all weights are equal, our bounds reduce to the bounds derived in the literature for the unweighted setting.
1 Introduction
The mathematical theory of fair item allocation among multiple agents has practical applications in scenarios like inheritance and partnership dissolutions. When agents have equal entitlements, as in inheritance cases, each agent naturally expects their allocated portion to be at least as good as others’. An allocation satisfying this requirement is called envy-free (EF).
When the items available for allocation are all indivisible, an envy-free allocation might not exist. A solution often applied in practice is to use money to compensate for the envy.
It is common to assume that the agents are quasilinear. This means that their utility equals their total value for the items they receive, plus the amount of money they receive (which may be positive or negative). Alternatively, one can consider a hypothetical third-party, who is willing to subsidize the process such that all agents receive a non-negative amount, and ask what is the minimum amount of subsidy that is required to attain envy-freeness.
The subsidy minimization problem was first studied by Halpern and Shah [15]. They showed that for any given allocation, there exists a permutation of the bundles that can be made envy-free with subsidies. The total required subsidy is at most $(n-1)mV$, where $m$ is the number of items, $n$ the number of agents, and $V$ the maximum item value for an agent, and this bound is tight in the worst case. Brustle et al. [7] considered the case in which the allocation is not given, but can be chosen. They presented an algorithm that finds an envy-freeable allocation through iterative maximum matching, requiring a total subsidy of at most $(n-1)V$, which is tight too.
This paper extends previous work on subsidy minimization by considering agents with different entitlements. In partnership dissolutions, partners often hold varying numbers of shares, entitling them to different proportions of the asset. Each agent expects to receive at least the same ”value per share” as others.
For example, if agent $i$ has twice the entitlement of agent $j$, $i$ expects a bundle worth at least twice as much as $j$’s.
Formally, an allocation is called weighted envy-free (WEF) (see e.g. [20, 32, 18]) if for every two agents $i$ and $j$, $\frac{1}{w_{i}}$ times the utility that $i$ assigns to his own bundle is at least as high as $\frac{1}{w_{j}}$ the utility that $i$ assigns to the bundle of $j$, where $w_{i}$ is $i$’s entitlement and $w_{j}$ is $j$’s entitlement. To illustrate the difficulty in this generalized setting, we show that both results mentioned above [15, 7] fail when agents have different entitlements.
Example 1.1.
There are two items $o_{1},o_{2}$ and two agents $i_{1},i_{2}$, with weights $w_{1}=1,w_{2}=10$ and valuation functions
$\begin{bmatrix}&o_{1}&o_{2}\\ i_{1}&5&7\\ i_{2}&10&8\\ \end{bmatrix}$
We will show that, contrary to the result of Halpern and Shah [15], there exists a division of items where no permutation satisfies WEF. Consider the bundles $A_{1}=\{o_{1}\}$ and $A_{2}=\{o_{2}\}$, where $i_{1}$ receives $A_{1}$ and $i_{2}$ receives $A_{2}$. Let $s_{1}$ and $s_{2}$ represent the subsidies for $i_{1}$ and $i_{2}$, respectively. The utility of $i_{1}$ for their own bundle is $5+s_{1}$, and for $i_{2}$’s bundle, it is $8+s_{2}$. To satisfy WEF, we need: $\frac{5+s_{1}}{1}\geq\frac{7+s_{2}}{10}$, which implies $s_{2}\leq 43+10s_{1}$. Similarly, for agent $i_{2}$, WEF requires: $\frac{8+s_{2}}{10}\geq\frac{10+s_{1}}{1}$, which implies $s_{2}\geq 92+10s_{1}$. These two conditions are contradictory, so no subsidies can make this allocation WEF. Next, consider the permutation where $i_{1}$ receives $A_{2}$ and $i_{2}$ receives $A_{1}$. In this case, WEF requires: $\frac{7+s_{1}}{1}\geq\frac{5+s_{2}}{10}$, which implies $s_{2}\leq 65+10s_{1}$. and for $i_{2}$: $\frac{10+s_{2}}{10}\geq\frac{8+s_{1}}{1}$, which implies $s_{2}\geq 70+10s_{1}$. Again, these conditions are contradictory, proving that no permutation of bundles satisfies WEF.
This example also shows that the Iterated Maximum Matching algorithm of Brustle et al. [7] does not guarantee WEF. The algorithm yields an allocation where all agents receive the same number of items, but as shown, no such allocation can be made WEF.
1.1 Related Work
Equal entitlements.
Steinhaus [22] initiated fair allocation with the cake-cutting problem, followed by Foley’s [13] advocacy for envy-free resource allocation. Challenges with indivisible items were outlined by Schmeidler and Yaari [21].
Binary valuations.
Particularly relevant to our work is a recent work by Viswanathan and Zick [25], who devised a fair allocation method inspired by Yankee Swap, achieving efficient and fair allocations when agents have submodular binary valuations. We use some of their techniques in our algorithms.
Subsidies.
The concept of compensating an indivisible resource allocation with money has been explored in the literature ever since Demange, Gale, and Sotomayor [12] introduced an ascending auction for envy-free allocation using monetary payments for unit demand agents.
In addition to their results for additive valuations mentioned earlier, Brustle et al. [7] also study the more general class of monotone valuations. They demonstrate that a total subsidy of $2(n-1)^{2}V$ is sufficient to guarantee the existence of an envy-freeable allocation. Kawase el al. [16] improved this bound to $\frac{n^{2}-n-1}{2}$.
Caragiannis and Ioannidis [8] developed an algorithm that approximates the minimum subsidies with any required accuracy for a constant number of agents, though with increased running time. However, for a super-constant number of agents, they showed that minimizing subsidies for envy-freeness is both hard to compute exactly and difficult to approximate.
Aziz [2] presented a sufficient condition and an algorithm to achieve envy-freeness and equitability (every agent should get the same utility) when monetary transfers are allowed for agents with quasi-linear utilities and superadditive valuations (positive or negative).
Barman et al. [6] studied agents with dichotomous valuations (agents whose marginal value for any good is either zero or one), without any additivity requirement. They proved that, for $n$ agents, there exists an allocation that achieves envy-freeness with total required subsidy of at most $n-1$, which is tight even for additive valuations.
Goko et al. [14] study an algorithm for an envy-free allocation with subsidy, that is also truthful, when agents have submodular binary valuations. The subsidy per agent is at most $V=1$. Their algorithm works only for agents with equal entitlements.
Different entitlements.
In the past few years, several researchers have examined a more general model in which different agents may have different entitlements, included weighted fairness models like weighted maximin share fairness (WMMS) and weighted proportionality up to one item (WPROP1) [9, 5, 3]. Chakraborty et al. [18] established maximum weighted Nash welfare (MWNW) satisfies Pareto optimality and introduced a weighted extension of EF1. Suksompong and Teh [23] demonstrated MWNW properties under binary additive valuations and its polynomial-time computability. They further extended these findings to various valuation types.
Different entitlements with subsidies.
Wu et al. ([31]) presented a polynomial-time algorithm for computing a PROP allocation of chores among agents with additive valuations, with total subsidy at most $\frac{nV}{4}$, which is tight. For agents with different entitlements, they compute a WPROP allocation with total subsidy at most $\frac{(n-1)V}{2}$. In a subsequent work ([30]), they further improved this bound to $(\frac{n}{3}-\frac{1}{6})V$.
As far as we know, weighted envy-freeness with subsides has not been studied yet. Our paper aims to fill this gap.
Matroid-rank valuations.
Recent studies have considered matroid-rank valuation (binary submodular). Montanari et al. [19] introduce a new family of weighted envy-freeness notions based on the concept of transferability and provide an algorithm for computing transferable allocations that maximize welfare. Babaioff et al. [4] design truthful allocation mechanisms that maximize welfare and are fair. Particularly relevant to our work is a recent work by Viswanathan and Zick [25], who devised a fair allocation method inspired by Yankee Swap, achieving efficient and fair allocations when agents have submodular binary valuations. We use some of their techniques in our algorithms. Later, Viswanathan and Zick in [26] generalize the Yankee Swap algorithm to efficiently compute allocations that maximize any fairness objective, called General Yankee Swap. We show that, under binary additive valuations, the general Yankee Swap algorithm can be adapted to agents with different entitlements. The adaptation is shown in Section 5.
1.2 Our Results
We derive bounds on the amount of subsidy required in order to attain a WEF allocation, in several different settings. We assume, without loss of generality, that the entitlements are ordered in increasing order, i.e., $w_{1}\leq w_{2}\leq\ldots\leq w_{n}$.
We denote $W:=\sum_{i=1}^{n}w_{i}$.
In Section 3.1, we assume that an allocation is given. As shown in Example 1.1, there are instances in which no rearrangement of bundles yields a WEF-able allocation. We prove a necessary and sufficient condition under which the allocation is WEF-able. We show that, when the allocation is WEF-able, a total subsidy of $(\frac{W}{w_{1}}-1)mV$ is sufficient to make it WEF, and prove that this bound is tight in the worst case.
This raises the question of whether a weighted-envy-free allocation with subsidy always exists? We provide an affirmative answer in Section 3.2. We show how to compute for agents with integer weights, a WEF-able allocation with total subsidy at most $\left(W-w_{1}\right)$.
Our algorithm extends Brustle et al. [7] algorithm, which in the unweighted setting attains total subsidy at most $(n-1)V$. With equal entitlements, we can normalize the weights to 1, maintaining the same bounds as in the unweighted case.
Following [15], in addition to the setting of general additive valuations, we study the setting in which agents have identical additive valuations (Section 4), and the setting in which agents have binary additive valuations (Section 5).
For identical additive valuations, we demonstrate an efficient computation of a WEF allocation with subsidies capped at $V$ for each agent, and a total subsidy bounded by $(n-1)V$, which is identical to the unweighted setting. We also prove that this bound is tight in the worst case.
Interestingly, in this special case (in contrast with the general case), the bound on the subsidy does not depend on the weights.
In Section 5, under binary-additive valuations, we show how to modify the General Yankee Swap algorithm of [26] to efficiently compute a WEF-able allocation with subsidies not exceeding $\frac{w_{i}}{w_{1}}$ per agent $i\in N$, and a total subsidy bounded by $\frac{W-w_{1}}{w_{1}}=\frac{W}{w_{1}}-1$, if all agents’ weights are integer. This bound reduces to an identical overall subsidy of $n-1$ in the unweighted setting.
Our findings and contributions are briefly summarized in Table  1.
Table 1: Distinctions between outcomes established in prior research (see citation), and those newly established in the present study, highlighted in bold.
All subsidy upper bounds are attainable by polynomial-time algorithms.
In the unweighted setting $w_{1}=\ldots=w_{n}=1$, and all upper bounds become $(n-1)V$.
Unweighted Setup Weighted Setup
General Valuations Identical Valuations Binary Valuations
Character- ization of WEF Allocation (1) No positive cost cycles, (2) USW maximi- zation [15] No positive-cost cycles. (3.2)
Permut- ation of a given allocation, that maximizes sum of values Always EF-able [15] Not necessarily WEF-able (3.3) Always WEF-able. (4.3) For non- redundant allocation: Always WEF-able (5.5)
Total subsidy upper bound $(n-1)V$ [15] $(W-w_{1})V$ (3.17) $(n-1)V$ (4.7) $\frac{W}{w_{1}}-1$ (5.8)
Subsidy bound for a given allocation $(n-1)mV$ [15] $\left(\frac{W}{w_{1}}-1\right)mV$ (3.5)
Total subsidy lower bound $(n-1)V$ [15] $\left(\frac{W}{w_{1}}-1\right)V$ (3.9) $(n-1)V$ (4.1) $\frac{W}{w_{2}}-1$ (5.1)
Remark 1.2.
Practical fair allocation cases use budget-balanced payments rather than subsidies. We use the subsidies terminology for consistency with previous works.
2 Preliminaries
Agents and valuations.
We denote by $[t]$ the set $\{1,2,...,t\}$ for any positive integer $t$. We focus on the problem of allocating $M=\{o_{1},...,o_{m}\}$ indivisible items among $N=[n]$ agents. Each subset of $M$ is called a bundle, and a combination of a bundle and a monetary transfer (positive or negative) is called a portion.
Each agent $i\in N$ has a valuation function $v_{i}:2^{M}\rightarrow\mathbb{R}_{\geq 0}$, indicating how much they value different bundles. For simplicity, we write $v_{i}(o_{1},...,o_{t})$ instead of $v_{i}(\{o_{1},...,o_{t}\})$. Additionally, we define $V=\displaystyle\max_{i\in N,o\in M}v_{i}(o)$.
The set $\Pi_{k}(S)$ denotes the collection of ordered partitions of a set $S\subseteq M$ into $k$ bundles. An allocation, denoted as $A\in\Pi_{n}(M)$, assigns the items to the agents. It consists of $n$ disjoint bundles $(A_{i})_{i\in N}$, where $\forall{i\in N}:A_{i}\subseteq M$, and $A_{i}\cap A_{j}=\emptyset$ for all $i\neq j\in N$. The bundle $A_{i}$ is given to agent $i\in N$, and $v_{i}(A_{i})$ represents agent $i$’s valuation of their bundle under allocation $A$. We consider only allocations $A$ that are complete, that is, $\cup_{i\in N}A_{i}=M$.
We assume that agents are quasilinear, so that the utility of each agent is $u_{i}(A_{i},s_{i}):=v_{i}(A_{i})+s_{i}$.
We assume that the valuation functions $v_{i}$ are normalized, that is $v_{i}(\emptyset)=0$ for all $i\in N$ (Suksompong and Teh. [24]); and additive, that is $v_{i}(S)=\sum_{o\in S}v_{i}(o)$ for all $S\subseteq M,i\in N$. In some sections of this paper, we make some additional assumptions:
  1. 1.
    identical additive: there exists an additive valuation function $v$ such that $v_{i}\equiv v$ for all $i\in N$.
  2. 2.
    binary additive: for all $i\in N$ and $o\in M$, $v_{i}(o)\in\{0,1\}$.
Without loss of generality, we assume that each item is valued positively by at least one agent; items that are valued at 00 by all agents can be allocated arbitrarily without affecting the envy.
Entitlements.
Each agent $i\in N$ is endowed with a fixed entitlement $w_{i}\in\mathbb{R}_{>0}$. We also refer to entitlement as weight. We assume, without loss of generality, that the entitlements are ordered in increasing order, i.e., $w_{1}\leq w_{2}\leq\ldots\leq w_{n}$. We denote $W:=\sum_{i\in N}w_{i}$.
Utilitarian social welfare.
The utilitarian social welfare (USW) of an allocation $A\in\Pi_{n}(M)$ is $USW(A)=\sum_{i\in N}v_{i}(A_{i})$. An allocation $A\in\Pi_{n}(M)$ is called maximizing utilitarian social welfare, denoted $MUSW^{M,N,v}$, if and only if $\sum_{i\in N}v_{i}(A_{i})\geq\sum_{i\in N}v_{i}(B_{i})$ for any other allocation $B\in\Pi_{n}(M)$.
The weighted utilitarian social welfare (WUSW) of an allocation $A\in\Pi_{n}(M)$ is the sum of values obtained by each agent, scaled by ratio of their weights, denoted as $WUSW(A)=\sum_{i\in N}w_{i}v_{i}(A_{i})$. An allocation $A\in\Pi_{n}(M)$ is considered $MWUSW^{M,N,v}$ if and only if $\sum_{i\in N}w_{i}v_{i}(A_{i})\geq\sum_{i\in N}w_{i}v_{i}(B_{i})$ for any other allocation $B\in\Pi_{n}(M)$.
To see why we multiply by the weights, consider a setting with one item and two agents, who value the item at $v_{1}=4$ and $v_{2}=6$. With equal weights, the MUSW allocation naturally gives the item to agent 2, who values it higher. However, if the weights are e.g. $w_{1}=3,w_{2}=1$, the MWUSW allocation gives the item to agent 1, since the weighted value $w_{1}v_{1}=12$ is larger than $w_{2}v_{2}=6$. Multiplying the values by the weights gives higher priority to the agent with the higher weight, as expected.
Envy.
The envy of agent $i$ towards agent $j$ under an allocation $A$ and subsidy vector $\mathbf{s}$ is defined as
$u_{i}(A_{j})-u_{i}(A_{i})=(v_{i}(A_{j})+s_{j})-(v_{i}(A_{i})+s_{i})$
If for all $i,j\in N$ the envy of $i$ towards $j$ is at most 00, then $(A,\mathbf{s})$ is called envy-free (EF).
The weighted-envy of agent $i$ towards agent $j$ under an allocation $A$ and subsidy $\mathbf{s}$ is defined as
$\frac{u_{i}(A_{j})}{w_{j}}-\frac{u_{i}(A_{i})}{w_{i}}=\frac{v_{i}(A_{j})+s_{j}}{w_{j}}-\frac{u_{i}(A_{i})+s_{i}}{w_{i}}$
If for all $i,j\in N$ the weighted envy of $i$ towards $j$ is at most 00, then $(A,\mathbf{s})$ is called a weighted-envy-free (WEF) solution.
To see why we divide by the weights, consider again a setting of one item and two agents who value it identically, $v_{1}=v_{2}=V$, and suppose the item is given to agent 1, and agent 2 gets some subsidy $s_{2}$. With equal entitlements, the envy of agent 2 is $(v_{2})-(s_{2})$, and a subsidy of $s_{2}=V$ is required to eliminate this envy. However, if the weights are $w_{1}=3,w_{2}=1$, the envy of agent 2 is $(v_{2}/w_{1})-(s_{2}/w_{2})$, and the required subsidy is only $s_{2}=V/3$; In contrast, if the weights are $w_{1}=1,w_{2}=3$, then the required subsidy is $s_{2}=3V$. This is expected, as there is less justification for envy when agent 1 has a higher entitlement, and more justification for envy when agent 2 has a higher entitlement.
Intuitively, the term $\frac{v_{i}(A_{i})}{w_{i}}$ represents the value per unit entitlement for agent $i$ in their allocation. The WEF condition ensures that this value is at least as high as $\frac{v_{i}(A_{j})}{w_{j}}$, denoting the corresponding value per unit entitlement for agent $j$ in the same allocation.
WEF seamlessly reduces to envy-free (EF) concept when entitlements are equal, i.e., $w_{i}=w_{j}$ for all $i,j\in N$. Like EF, which can be challenging to achieve with indivisible items, WEF faces similar limitations in guaranteeing fairness under such circumstances.
Often, we are given only an allocation $A$, and want to find a subsidy vector with which the allocation will be WEF.
Definition 2.1.
(a) We say that subsidy vector $\mathbf{s}$ is weighted-envy-eliminating for allocation $A$ if $(A,\mathbf{s})$ is WEF.
(b) An allocation $(A_{i})_{i\in N}$ is called weighted-envy-freeable (WEF-able) if there exists a subsidies vector $\mathbf{s}$ such that $(A,\mathbf{s})$ is WEF.
In the setting without money, WEF can be relaxed by allowing envy up to an upper bound that depends on the item values. We employ the generalization of allowable envy proposed by Chakraborty et al. [10]:
Definition 2.2.
(Chakraborty et al.[10]). For $x,y\in[0,1]$, an allocation $A$ is said to satisfy WEF(x,y) if for any $i,j\in N$, there exists $B\subseteq A_{j}$ with $|B|\leq 1$ such that $\frac{v_{i}(A_{i})+yv_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j})-xv_{i}(B)}{w_{j}}$, where $x+y=1$.
Specifically, WEF(0,0) corresponds to WEF.
Weighted envy graph.
The weighted envy graph of allocation $A$, denoted $G_{A,w}$, is a complete directed graph consisting of a set of vertices representing agents $N$, each assigned a weight denoted by $w_{i}$. In case of identical weights, we denote it $G_{A}$.
For any pair of agents $i$ and $j$ in $N$, the cost assigned to the arc $(i,j)$ in $G_{A,w}$ is defined as the weighted envy that agent $i$ holds toward agent $j$ under allocation $A$:
$cost_{A}(i,j):=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}$
We denote the cost of a path $(i_{1},...,i_{k})$ as $cost_{A}(i_{1},...,i_{k})=\sum_{j=1}^{k-1}cost_{A}(i_{j},i_{j+1})$. With these definitions, $\ell_{i,j}(A)$ represents the maximum-cost path from $i$ to $j$ in $G_{A,w}$, and $\ell_{i}(A)$ represents the maximum-cost path in $G_{A,w}$ starting at $i$.
The previous work of Halpern and Shah [15] provides sufficient and necessary conditions for an EF-able allocation in the unweighted setup. The following theorem considers a pre-determined allocation $A=(A_{1},\ldots,A_{n})$, where we aim to assign one bundle to each agent.
Theorem 2.3.
(Halpern and Shah [15]). In a setting with equal entitlements, the following statements are equivalent:
  1. 1.
    The allocation $A$ is envy-freeable.
  2. 2.
    The allocation $A$ maximizes the utilitarian welfare across all reassignments of its bundles to agents, that is, for every permutation $\sigma$ of $[n]$ (a bijection $\sum:[n]\rightarrow[n]$), $\sum_{i\in N}v_{i}(A_{i})\geq\sum_{i\in N}v_{i}(A_{\sigma{i}})$.
  3. 3.
    $G_{A}$ has no positive-cost cycles.
The theorem implies that, in the unweighted setup, for every allocation there exists a reassignment of bundles allocated to the agents, ensuring its envy-freeability.
Halpern and Shah [15] have outlined a method to find a subsidy vector $\mathbf{s}$, ensuring envy-freeness while minimizing the required subsidy. Each agent’s subsidy is determined by the maximum cost path from that agent in the envy graph. This cost can be computed within strongly polynomial time.
Furthermore, they identify the subsidy needed in the worst case for two scenarios:
  1. 1.
    When the allocation is given: the minimum subsidy required is $(n-1)mV$ in the worst case. [15].
  2. 2.
    When the allocation can be chosen: the minimum subsidy required is at least $(n-1)V$ in the worst case, even in the special case of binary valuations [15].
3 WEF-able Allocations
In this section, we expand the idea of fairness by considering weights and discuss the challenge of identifying WEF-able allocations, compared to EF-able allocations.
We demonstrate how to compute a subsidy vector that eliminates envy among the agents for a given allocation. Naturally, the question arises whether a WEF-able allocation for agents with general additive valuations always exists. We address and answer this inquiry.
3.1 A given allocation: the weighted-envy graph
When trying to extend Theorem 2.3 to the weighted setup, a notable distinction arises: in the unweighted context, an allocation $A$ is EF-able if and only if $A$ maximizes the utilitarian welfare across all reassignments of its bundles to agents (part (2) of Theorem 2.3).
Formally, for every permutation $\sigma$ of $[n]$, it holds that
$\sum_{i\in N}(v_{i}(A_{i}))\geq\sum_{i\in N}(v_{i}(A_{\sigma(i)}))$.
Contrastingly, in the weighted setup, this assertion does not necessarily hold, as demonstrated by Example 1.1. Moreover, the converse of (3) in Theorem 2.3 can fail too.
Proposition 3.1.
There exists an allocation $A$, which is WEF-able, yet maximizes neither the utilitarian social welfare or the weighted utilitarian social welfare.
Proof.
Consider two agents with weights $w_{1}=2,w_{2}=3$ and $2$ items. The valuation functions are as follows:
$\begin{bmatrix}&o_{1}&o_{2}\\ i_{1}&8&10\\ i_{2}&6&7\\ \end{bmatrix}$
Consider the allocation $A=(A_{1},A_{2})$ when $A_{1}=\{o_{1}\},A_{2}=\{o_{2}\}$, and subsidies $s_{1}=0,s_{2}=2$:
  1. 1.
    The envy of agent $i_{1}$ towards agent $i_{2}$ is $\frac{v_{1}(A_{2})+s_{2}}{w_{2}}-\frac{v_{1}(A_{1})+s_{1}}{w_{1}}=\frac{10+2}{3}-\frac{8+0}{2}=0$
  2. 2.
    the envy of agent $i_{2}$ towards agent $i_{1}$ is $\frac{v_{2}(A_{1})+s_{1}}{w_{1}}-\frac{v_{2}(A_{2})+s_{2}}{w_{2}}=\frac{6+0}{2}-\frac{7+2}{3}=0$
so the allocation is WEF. However, we can achieve both higher USW and higher WUSW by swapping the bundles: $A_{1}^{{}^{\prime}}=A_{2},A_{2}^{{}^{\prime}}=A_{1}$, as
$v_{1}(A_{1})+v_{2}(A_{2})=8+7=15<$
$16=10+6=v_{1}(A_{1}^{\prime})+v_{2}(A_{2}^{\prime})$
$w_{1}v_{1}(A_{1})+w_{2}v_{2}(A_{2})=16+21=37<$
$38=20+18=w_{1}v_{1}(A_{1}^{\prime})+w_{2}v_{2}(A_{2}^{\prime}).\qed$
It turns out that two out of three parts of this theorem still hold in the weighted setting; The proof is similar to the one in [15].
Theorem 3.2.
For an allocation $A$, the following statements are equivalent:
  1. (a)
    $A$ is weighted-envy-freeable
  2. (b)
    $G_{A,w}$ has no positive-cost cycles.
Proof.
We show $(a)\Rightarrow(b)$ and $(b)\Rightarrow(a)$.
For $(a)\Rightarrow(b)$, we prove a more general claim: the total cost of any cycle in $G_{A,w}$ is the same for any subsidy vector. Indeed, suppose we give some agent $i$ a subsidy of $s_{i}$. As a result, the cost of every edge from $i$ decreases by $s_{i}/w_{i}$ (as $i$ experiences less envy), and the cost of every edge into $i$ increases by $s_{i}/w_{i}$ (as other agents experience more envy in $i$). Every cycle through $i$ contains exactly one edge from $i$ and one edge into $i$, and every other cycle contains no such edges. Therefore, the total cost of any cycle remains unchanged.
Now, if $A$ is WEF-able, then for some subsidy vector, the costs of all edges are at most 00, so the total costs of all cycles are at most 00. Therefore, the costs of all cycles are at most 00 even without the subsidy vector.
For $(b)\Rightarrow(a)$, we present a specific subsidy vector. Suppose $G_{A,w}$ has no positive-cost cycles. Then, $\ell_{i}(A)$ (the maximum cost of a path in $G_{A,w}$ starting at $i$) is well-defined and finite. For each $i\in N$, let $s_{i}=w_{i}cost_{A}(\ell_{i}(A))$. It is noteworthy that $s_{i}\geq 0$, since there is a path from $i$ to $i$ with 00 cost, establishing the suitability of s as a valid subsidies vector.
Furthermore, the following holds for all pairs $i\neq j\in N$:
$\frac{s_{i}}{w_{i}}$ $=cost_{A}(\ell_{i}(A))\geq cost_{A}(i,j)+cost_{A}(\ell_{j}(A))$
$=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}+cost_{A}(\ell_{j}(A))$
$=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}+\frac{s_{j}}{w_{j}}$
This implies $\frac{v_{i}(A_{i})+s_{i}}{w_{i}}\geq\frac{v_{i}(A_{j})+s_{j}}{w_{j}}$. Hence, $(A,\mathbf{s})$ is weighted-envy-free, and thus, $A$ is weighted-envy-freeable. ∎
Theorem 3.2 presents an effective method for verifying whether a given allocation $A$ is WEF-able. We use it to present an alternative proof that part 2 of Theorem 2.3 does not hold in the weighted setting.
Proposition 3.3.
There exist bundles $B_{1},...,B_{n}$ such that for every permutation of the agents $\sigma:[n]\rightarrow[n]$, the resulting allocation $A_{i}=B_{\sigma(i)}$ is not WEF-able.
Proof.
As in Example 1.1, consider 2 bundles $B_{1},B_{2}$ and 2 agents, $N=\{i_{1},i_{2}\}$, with weights $w_{1}=1,w_{2}=10$ and valuation functions
$\begin{bmatrix}&B_{1}&B_{2}\\ i_{1}&5&7\\ i_{2}&10&8\\ \end{bmatrix}$
We show that for every permutation of the agents $\sigma:[n]\rightarrow[n]$, the resulting allocation $A_{i}=B_{\sigma(i)}$ results in the corresponding weighted envy graph a positive cost cycle.
The allocation $A_{1}=B_{1},A_{2}=B_{2}$ maximizes both the utilitarian social welfare and the weighted utilitarian social welfare. The weighted envy of agent $i_{1}$ towards agent $i_{2}$ is $cost_{A}(i_{1},i_{2})=\frac{v_{i_{1}}(A_{i_{2}})}{w_{i_{2}}}-\frac{v_{i_{1}}(A_{i_{1}})}{w_{i_{1}}}=\frac{7}{10}-5=-\frac{43}{10}$; the weighted envy of agent $i_{2}$ towards agent $i_{1}$ is calculated as $cost_{A}(i_{2},i_{1})=\frac{v_{i_{2}}(A_{i_{1}})}{w_{i_{1}}}-\frac{v_{i_{2}}(A_{i_{2}})}{w_{i_{2}}}=10-\frac{8}{10}=\frac{92}{10}$. Thus, the cost of the cycle $(i_{1},i_{2})$ is $\frac{49}{10}>0$, indicating a positive-cost cycle.
Also, if an attempt is made to swap the bundles in a circular manner, such that $A_{1}=A_{2},A_{2}=B_{1}$, the cost of this cycle $(i_{1},i_{2})$ is $(\frac{5}{10}-7)+(8-\frac{10}{10})=\frac{1}{2}>0$, affirming a positive cost as well.
By Proposition 3.2, both permutations are not WEF-able. ∎
Deciding that an allocation $A$ is WEF-able is not enough. As in Halpern and Shah [15], we also need to find a minimal subsidy vector, $\mathbf{s}$, that ensures WEF for $(A,\mathbf{s})$. This subsidy vector can be computed in strongly polynomial time and matches the one constructed in the proof of Proposition 3.2.
Theorem 3.4.
For any allocation $A$ that is WEF-able, let $\mathbf{s}^{*}$ be a subsidy vector defined by $\mathbf{s}^{*}_{i}:=w_{i}cost_{A}(\ell_{i}(A))$, for all $i\in N$. Then
  1. 1.
    $(A,\mathbf{s}^{*})$ is WEF;
  2. 2.
    Any other envy-eliminating subsidy vector $\mathbf{s}$ satisfies $s^{*}_{i}\leq s_{i}$ for all $i\in N$;
  3. 3.
    The computation of $s^{*}$ can be performed in $O(nm+n^{3})$ time.
  4. 4.
    There exists at least one agent $i\in N$ for whom $s^{*}_{i}=0$.
Proof.
  1. 1.
    The establishment of condition (b) implying condition (a) in Theorem 3.2 has already demonstrated the inclusion of $(A,\mathbf{s}^{*})$ is WEF.
  2. 2.
    Let $\mathbf{s}$ be a subsidy vector, such that $(A,\mathbf{s})$ is WEF, and $i\in N$ be fixed. Consider the highest-cost path originating from $i$ in the graph $G_{A,w}$. This path is denoted as $(i_{1},...,i_{k})$, with $i_{1}=i$ and $cost_{A}(i_{1},...,i_{r})=cost_{A}(\ell_{i}(A))=\frac{s^{*}_{i}}{w_{i}}$.
    Due to the WEF nature of $(A,\mathbf{s})$, it follows that for each $k\in[r-1]$, the following inequality holds:
    $\frac{v_{i_{k}}(A_{i_{k}})+s_{i_{k}}}{w_{i_{k}}}\geq\frac{v_{i_{k}}(A_{i_{k+1}})+s_{i_{k+1}}}{w_{i_{k+1}}}$
    $\Rightarrow\frac{s_{i_{k}}}{w_{i_{k}}}-\frac{s_{i_{k+1}}}{w_{i_{k+1}}}\geq\frac{v_{i_{k}}(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}=$
    $cost_{A}(i_{k},i_{k+1})$
    Summing this inequality over all $k\in[r-1]$, the following relation is obtained:
    $\frac{s_{i}}{w_{i}}-\frac{s_{i_{r}}}{w_{i_{r}}}=\frac{s_{i_{1}}}{w_{i_{1}}}-\frac{s_{i_{r}}}{w_{i_{r}}}\geq cost_{A}(i_{1},...,i_{r})=\frac{s_{i}^{*}}{w_{i}}$
    $\Rightarrow\frac{s_{i}}{w_{i}}\geq\frac{s_{i}^{*}}{w_{i}}+\frac{s_{i_{r}}}{w_{i_{r}}}\geq\frac{s_{i}^{*}}{w_{i}}$
    The final transition is valid due to the non-negativity of subsidies and weights, that is, $\frac{s_{i_{r}}}{w_{i_{r}}}\geq 0$.
  3. 3.
    The computation of $s^{*}$ can be executed as follows: Initially, the Floyd-Marshall algorithm ([28], [29]) is applied to the graph derived by negating all edge costs in $G_{A,w}$ (This has a linear time solution since there are no cycles with positive costs in the graph). Hence, determining the longest path cost between any two agents, accomplished in $O(nm+n^{3})$ time. Subsequently, the longest path starting at each agent is identified in $O(n^{2})$ time.
  4. 4.
    Assume, for the sake of contradiction, that $s^{*}_{i}>0$ for every agent $i\in N$, which implies that $\text{cost}_{A}(\ell_{i}(A))>0$ since $w_{i}\geq 1$. Because the number of agents is finite, there must be $k\geq 2$ agents, say $i_{1},\dots,i_{k}$, such that $\ell_{i_{1}}(A),\dots,\ell_{i_{k}}(A)$ form a cycle with positive cost in $G_{w,A}$, contradicting Theorem 3.2.
    Therefore, there must be at least one agent whose subsidy is 00.
Now we can find the minimum subsidy needed in the worst-case scenario for agents with different entitlements, whether the allocation is given or can be chosen. The proofs extend those of Halpern and Shah [15]. Note that Theorem 3.5 applies to any additive valuations, including identical and binary cases.
Theorem 3.5.
For every weight vector, and for every given WEF-able allocation $A$, there exists an envy-eliminating subsidy vector with total subsidy at most $\left(\frac{W}{w_{1}}-1\right)mV$, and this bound is tight in the worst case.
Proof.
For the lower bound, consider an instance with $m$ identical items, which all $n$ agents value at $V$. Consider the allocation $A$, which assigns all items to a single agent $j$ with the minimum entitlement. It’s evident that $A$ is WEF-able, and its optimal subsidy vector $\mathbf{s}$ satisfies $s_{j}=0$ and $s_{i}=\frac{w_{i}}{w_{j}}mV=\frac{w_{i}}{w_{1}}mV$ for $i\neq j\in N$. Therefore, we require
$\frac{W-w_{1}}{w_{1}}mV=\left(\frac{W}{w_{1}}-1\right)mV$
To establish the upper bound, the required subsidy amount is the summation of the maximum cost paths originating from different agents (Theorem 3.4). To bound the highest cost of a path starting from $i$, we make the following lemma:
Lemma 3.6.
For every WEF-able allocation $A$ and agent $i$, the highest cost of a path from $i$ in $G_{A,\mathbf{w}}$ is at most $\frac{mV}{w_{1}}$.
By Lemma 3.6, the highest cost of a path from $i$ in $G_{A,\mathbf{w}}$ is at most $\frac{mV}{w_{1}}$. So the subsidy of agent $i$ does not exceed $w_{i}\frac{mV}{w_{1}}$. By part (4) of 3.4, at least one agent has a subsidy of 0. This implies a total subsidy of at most
$\frac{W-w_{1}}{w_{1}}mV=\left(\frac{W}{w_{1}}-1\right)mV\qed$
Proof.
(Lemma 3.6) Since $A$ is WEF-able, there are no positive-weight cycles in $G_{A,w}$, so the highest cost of a path is finite.
For every two agents $i,j$,
$cost_{A}(i,j)=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq\frac{\sum_{o\in A_{j}}v_{i}(o)}{w_{1}}\leq\frac{\sum_{o\in M}v_{i}(o)}{w_{1}}$
Therefore, the cost of every path is at most $\frac{mV}{w_{1}}$. ∎
Importantly, for a fixed allocation, the weighted setting maximum worst-case subsidy depends on the proportion of $w_{1}$ in the total weight $W$, which can be much larger than the number of agents $n$, the argument that the unweighted setting depends on.
Halpern and Shah proved in [15] that, if an allocation $A$ is EF-able and EF1, then every path in $G_{A,w}$ has cost at most $(n-1)V$, and therefore a subsidy of at most $(n-1)^{2}V$ is sufficient. The following theorem generalizes this result to the weighted setting.
Theorem 3.7.
Let $A$ be both WEF-able and $WEF(x,y)$ for any $x,y\in[0,1]$ such that $x+y\leq 1$. Then there exists an envy-eliminating subsidy vector with total subsidy at most $\left(\frac{W}{w_{1}}-1\right)(n-1)V$.
Proof.
$A$ is $WEF(x,y)$, so for all $i,j\in N$, there exists $B\subseteq A_{j}$ with $|B|\leq 1$ where
$\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq\frac{yv_{i}(B)}{w_{i}}+\frac{xv_{i}(B)}{w_{j}}=\frac{(1-x)v_{i}(B)}{w_{i}}+\frac{xv_{i}(B)}{w_{j}}\leq$
$\frac{(1-x)v_{i}(B)}{w_{1}}+\frac{xv_{i}(B)}{w_{1}}=\frac{(1-x)v_{i}(B)+xv_{i}(B)}{w_{1}}=\frac{v_{i}(B)}{w_{1}}\leq\frac{V}{w_{1}}$
Any path contains at most $n-1$ arcs, that is, $s_{i}=w_{i}cost_{A}(\ell_{i}(A))\leq\frac{w_{i}}{w_{1}}(n-1)V$. From part (4) of Theorem 3.4, there is at least one agent that requires no subsidy, so the required total subsidy is at most $\frac{W-w_{1}}{w_{1}}(n-1)V=\left(\frac{W}{w_{1}}-1\right)(n-1)V$.
3.2 Computing a WEF-able allocation
We start with a lower bound on the required subsidy for general additive valuations. To prove the lower bound, we need a lemma.
Lemma 3.8.
Suppose there are $n$ agents, and only one item $o$ which the agents value positively. Then an allocation is WEF-able iff $o$ is given to an agent $i$ with the highest $v_{i}(o)$.
Proof.
By Theorem 3.2, it is sufficient to check the cycles in the weighted envy-graph. If $o$ is given to $i$, then $i$’s envy is $-\frac{v_{i}(o)}{w_{i}}$, and the envy of every other agent $j$ in $i$ is $\frac{v_{j}(o)}{w_{i}}$. All other envies are 0. The only potential positive-weight cycles are cycles of length 2 involving agent $i$. The weight of such a cycle is positive iff $\frac{v_{j}(o)}{w_{i}}-\frac{v_{i}(o)}{w_{i}}>0$, which holds iff $v_{j}(o)>v_{i}(o)$. Therefore, there are no positive-weight cycles iff $v_{i}(o)$ is maximum. ∎
Theorem 3.9.
For every weight vector $\mathbf{w}$ and any $n\geq 2$, no algorithm can guarantee a total subsidy smaller than $(\frac{W}{w_{1}}-1)V$.
Proof.
Consider an instance with $n$ agents with weights $w_{1}=1$ and $w_{i}>1$ for $i\geq 2$. There is one item $o$ with valuations $v_{1}(o)=V$ and $v_{i}(o)=V-\epsilon=V_{2nd}$ for $i\geq 2$, and the other items are worth 00 to all agents.
By Lemma 3.8, the only WEF-able allocation is to give $o$ to agent $1$. In this case, the minimum subsidy is $s_{1}=0$ and $s_{i}=w_{i}\frac{v_{i}(o)}{w_{1}}=w_{i}V_{2nd}$. Summing all subsidies leads to $\sum_{i\geq 2}w_{i}V_{2nd}=(W-1)V_{2nd}$. As $\epsilon$ can be arbitrarily small, we get a lower bound of $(W-1)V$. ∎
Our task is to construct a WEF-able allocation $A$ with maximum subsidy of $w_{i}V$ per agent $i\in N$. In other words, the maximum cost of a path needs to be below $V$.
For the unweighted setting, Brustle et al. [7] presented an algorithm, based on iterative maximum matching between agents and items, that finds an envy-freeable allocation for which the required subsidy is at most $(n-1)V$, which is tight.
However, this algorithm might result in an allocation that is not WEF-able, as demonstrated in Example 1.1: The allocation $A_{1}=\{o_{1}\},A_{2}=\{o_{2}\}$ is derived from maximum matching between the agents and items but is not WEF-able. Intuitively, since agent 1 has greater entitlements, he should receive a more valuable bundle.
Our algorithm addresses this by finding a maximum matching between several copies of the agents, based on their entitlements, and the items.
Since our algorithm operates with copies of agents according to their entitlements, we assume the entitlements are integers. In other words, the algorithm constructs a WEF-compatible allocation for agents with integer weights.
For simplicity, if the number of items at the start of a round is less than $W$, we add dummy items (valued at 00 by all agents) such that the total number of items is at least $W$. This implies that, in each round, every agent $i$ gets exactly $w_{i}$ items (each agent can receive at most $w_{i}$ items since they have $w_{i}$ copies. Because the algorithm computes a maximum matching between the $W=\Sigma_{i\in N}w_{i}$ items and the agents’ copies, each agent must receive $w_{i}$ items).
In example 1.1, we add 9 dummy items, and then find a maximum-weight matching between $w_{1}=10$ copies of agent 1 and $w_{2}=1$ copy of agent 2 with the items. The first round of maximum weight matching gives a real item to two copies of agent 1, resulting in the allocation $A_{1}=\{o_{1},o_{2}\},A_{2}=\emptyset$, which is WEF-able.
We denote by $H[I,J]$ the complete bipartite graph on vertex sets $I$ and $J$, where edge $(i,o_{j})$ has the value $v_{i}(o_{j})$.
In round $t$, the algorithm finds a maximum-value matching $\mu^{t}$ in $H[I,O_{t}]$, where $I$ is the set of copies of the agents and $O_{t}$ consists of the items that have not yet been allocated in this round. If the $w_{i}$ copies of agent $i$, denoted $i_{1},...,i_{w_{i}}$ are matched to items $\mu_{i_{1}}^{t},...,\mu_{i_{w_{i}}}^{t}$, respectively, then the algorithm allocates these items to agent $i$. We denote the total set of items allocated to $i$ as $A_{i}^{t}=\{\mu_{i_{1}}^{t},...,\mu_{i_{w_{i}}}^{t}\}$ and say that the algorithm computed the allocation $A^{t}$ at iteration $t$.
This process continues until iteration $T$, at which all items have been allocated.
Algorithm 1 Weighted Sequence Protocol For Additive Valuations
1:Instance $(N,M,v,\{w_{i}\}_{i\in N})$ with additive valuations.
2:WEF-able allocation $A=(A_{i})_{i\in N}$ with total required subsidy of at most $(W-1)V$.
3:$A_{i}\leftarrow\emptyset,\forall i\in N$
4:$t\leftarrow 1$; $O_{1}\leftarrow M$
[…]
6:while $O_{t}\neq\emptyset$ do
7:     Compute a maximum-value matching $\mu^{t}=\{(i,\mu_{i}^{t})\}_{i\in I}$ in $H[I,O_{t}]$
8:     $A_{i}^{t}\leftarrow\{\mu_{i_{1}}^{t},...,\mu_{i_{w_{i}}}^{t}\}$, for all $i\in N$
9:     Set $A_{i}\leftarrow A_{i}\cup A_{i}^{t}$, for all $i\in N$
[…]
11:     $t\leftarrow t+1$
12:end while
13:return $A$
First, we prove that the allocation $A$ output from Algorithm 1 is WEF-able. By Theorem 3.2, it is sufficient to prove that its weighted envy-graph has no positive-cost cycles.
We prove that this holds not only for the final allocation, but also for each allocation $A^{t}$ computed in iteration $t$.
For each cycle $C$ in the envy-graph of $A^{t}$, we construct another allocation by choosing one item uniformly at random from the bundle of each agent in $C$, and transferring this item to the following agent in $C$.
We show that the expected change in value between the new allocation and the allocation returned by Algorithm 1 is at most 00; we further show that the expected change is equal to the cost of $C$. Therefore, the cost of $C$ is at most 00.
Proposition 3.10.
For each round $t$ in Algorithm 1, $A^{t}$ is WEF-able.
Proof.
We prove that, in every round $t$, the total cost added to any directed cycle in the weighted-envy graph is non-positive. Combined with Theorem 3.2, this shows that $A^{t}$ is WEF-able for every round $t\in T$.
Let $A^{t}$ the allocation computed by Algorithm 1 at iteration $t$. Let $C$ be any directed cycle in $G_{A^{t},w}$, and denote $C=(i_{1},...,i_{r})$. To simplify notation, we consider $i_{1}$ as $i_{r+1}$.
We construct another allocation $B^{t}$ randomly as follows: for each agent $i_{j}\in C$, we choose one item $o_{i_{j+1}}^{t}$ uniformly from $i_{j+1}$’s bundle and transfer it to $i_{j}$’s bundle  [^1] Recall that at each iteration, each agent $i_{j}$ receives exactly $w_{i_{j}}$ items.. The expected value of $v_{i_{j}}(o_{i_{j+1}}^{t})$, the value of the item removed from $i_{j}$’s bundle, can be computed as the average value of all items in $A^{t}_{i_{j}}$: $\frac{\sum_{o\in A^{t}_{i_{j}}}v_{i_{j}}(o)}{w_{i_{j}}}=\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}$. Similarly, the expected value of $v_{i_{j}}(o_{i_{j}}^{t})$, the value of the item added the $i_{j}$’s bundle, is $\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}$. Thus, the expected change in value between $B^{t}$ and $A^{t}$ is
$\mathbb{E}\left[\sum_{i\in N}\left(v_{i}(B_{i}^{t})-v_{i}(A^{t}_{i})\right)\right]=\sum_{i_{j}\in C}\left(\mathbb{E}\left[v_{i_{j}}(o_{i_{j+1}}^{t})\right]-\mathbb{E}\left[v_{i_{j}}(o_{i_{j}}^{t})\right]\right)=$
$\sum_{i_{j}\in C}\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}$
This is exactly the total cost of cycle $C$. According to Algorithm 1, $A^{t}$ maximizes the total value among all allocations in which each agent $i$ receives exactly $w_{i}$ items. Therefore, the left-hand side of the above expression must be at most 00. But the right-hand side of the same expression is exactly the total cost of $C$. Therefore,
$0\geq\mathbb{E}\left[\sum_{i\in N}\left(v_{i}(B^{t}_{i})-v_{i}(A^{t}_{i})\right)\right]=cost_{A^{t}}(C),$
so the cost of every directed cycle is at most 00, as required. ∎
Next, we show that the weighted envy between agents is at most $V$.
Proposition 3.11.
Let $A$ be the allocation computed by Algorithm 1. For every $i,j\in N$ it holds that $\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq V$
Proof.
For $t\in[T]$, Algorithm 1 computes a maximum-value matching $\mu^{t}$ in $H[I,O_{t}]$ and allocates the items $\mu_{i_{1}}^{t}\ldots\mu_{i_{w_{i}}}^{t}$ to agent $i$. By maximality of the matching, the items assigned in round $t$ are at least as valuable as the items remaining after that round, that is, $v_{i}(o_{1})\geq v_{i}(o_{2})$ for each $i\in N$, $o_{1}\in A_{i}^{t}$ and $o_{2}\in O_{t+1}$. Thus, for each round $t\in[T]$, for $i\neq j\in N$ and $o\in O_{t+1}$ that $i$ values the most, it holds that
$\frac{v_{i}(A_{i}^{t})}{w_{i}}=\frac{v_{i}(\{\mu_{i_{1}}^{t}\cup\ldots\cup\mu_{i_{w_{i}}}^{t}\})}{w_{i}}\geq\frac{w_{i}v_{i}(o)}{w_{i}}=$
$\frac{w_{j}v_{i}(o)}{w_{j}}\geq\frac{v_{i}(\{\mu_{j_{1}}^{t+1}\cup\ldots\cup\mu_{j_{w_{j}}}^{t+1}\})}{w_{j}}=\frac{v_{i}(A_{j}^{t+1})}{w_{j}}$
Therefore,
$\frac{v_{i}(A_{i})}{w_{i}}=\sum_{t=1}^{T}\frac{v_{i}(A_{i}^{t})}{w_{i}}\geq\sum_{t=1}^{T-1}\frac{v_{i}(A_{i}^{t})}{w_{i}}\geq\sum_{t=2}^{T}\frac{v_{i}(A_{j}^{t})}{w_{j}}=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{j}^{1})}{w_{j}}$
Hence, the envy of agent $i$ for agent $j$ is bounded by
$\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq\frac{v_{i}(A_{j}^{1})}{w_{j}}\leq\frac{w_{j}V}{w_{j}}=V.\qed$
Proposition 3.11 implies an upper bound of $(n-1)(W-w_{1})V$ on the total subsidy, but we can show an even better bound by bounding the subsidy per agent. First we generalize Lemma 4.1 of Brustle et al. [7] to the weighted setting:
Lemma 3.12.
Let $A$ be a WEF-able allocation. For any positive number $v$, If $cost_{A}(i,k)\geq-v$ for every arc $(i,k)$ in $G_{A,w}$, then the maximum subsidy required is at most $w_{i}v$ per agent $i\in N$.
Proof.
Assume $\ell_{i}(A)=(i\ldots j)$.
Then it holds for the cycle $C=(i\ldots j)$ that
$cost_{A}(C)=cost_{A}(\ell_{i}(A))+cost_{A}(j,i).$
By Theorem 3.2, $cost_{A}(C)\leq 0$, thus,
$cost_{A}(\ell_{i}(A))\leq-cost_{A}(j,i)\leq V.$
Therefore, $s_{i}=w_{i}cost_{A}(\ell_{i}(A))\leq w_{i}V$. ∎
In most allocations, including the one resulting from Algorithm 1, it is not always true that $cost_{A}(i,k)\geq-V$ for every arc $(i,k)$. We apply the technique from Brustle et al. [7] to use Lemma 3.12 with a modified valuation function $\bar{v}_{i}$, derived from the weighted valuation $\frac{v_{i}(A_{i})}{w_{i}}$.
Let $A^{t}$ be the output allocation from Algorithm 1, computed in iteration $t$. For each $i\in N$, we define the modified valuation function as follow:
[…]
Under the modified valuations, for any two agents $i,j\in N$, the modified-cost assigned to the arc $(i,j)$ in the envy graph (with unit weights) is defined as $\overline{cost}_{A}(i,j)=\bar{v}_{i}(A_{i})-\bar{v}_{i}(A_{j})$. Moreover, the modified-cost of a path $(i_{1},...,i_{k})$ is $\overline{cost}_{A}(i_{1},...,i_{k})=\sum_{j=1}^{k-1}\overline{cost}_{A}(i_{j},i_{j+1})$.
We will prove that an allocation that is WEF-able for the original valuations is also WEF-able for the modified valuations (Proposition 3.14), and that the maximum subsidy required by each agent for the original valuations is bounded by the subsidy required for the modified valuations (Proposition 3.15).
Next, we demonstrate that under the modified valuations, the cost of each arc is at least $-V$. Finally, by Lemma 3.12, we conclude that the maximum subsidy required for any agent $i\in N$ is $w_{i}V$ for the modified valuation $\bar{v}$ (3.16) and for the original valuations $v$ as well.
Observation 3.13.
For agent $i\in N$ and round $t\in[T]$ it holds that:
  1. 1.
    $\bar{v}_{i}(A_{i}^{t})=\frac{v_{i}(A_{i}^{t})}{w_{i}}$.
  2. 2.
    For agent $j\neq i\in N$, $\bar{v}_{i}(A_{j}^{t})\geq\frac{v_{i}(A_{j}^{t})}{w_{j}}$; hence $\bar{v}_{i}(A_{j}^{t})-\bar{v}_{i}(A_{i}^{t})\geq\frac{v_{i}(A_{j}^{t})}{w_{j}}-\frac{v_{i}(A_{i}^{t})}{w_{i}}$.
Proposition 3.14.
Assume $A$ is WEF-able under the original valuations $v$. Then, $A$ is EF-able (i.e., WEF-able with unit weights) under the modified valuations $\bar{v}$.
Proof.
By Theorem 3.2, it is sufficient to prove that all directed cycles in the envy graph (with the modified valuations and unit weights) have non-positive total cost. We prove a stronger claim: in every round $t$, the total modified-cost added to every directed cycle $C$ is non-positive. Let $A^{t}$ be the allocation computed by Algorithm 1 at iteration $t$. Suppose, contrary to our assumption, that there exists a cycle $C=(i_{1},...,i_{r})$ and a round $t$ in which the modified-cost added to $C$ is positive. To simplify notation, we consider $i_{1}$ as $i_{r+1}$. This implies that
$\sum_{j=1}^{r}\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})>\sum_{j=1}^{r}\bar{v}_{i_{j}}(A_{i_{j}}^{t}).$ (1)
There are several cases to consider.
Case 1: All arcs $i_{j}\to i_{j+1}$ in $C$ have $\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})=\frac{v_{i_{j}}(A_{i_{j+1}}^{t})}{w_{i_{j+1}}}$ (in particular, this holds for $t=T$). In this case, inequality (1) implies
$\overline{cost}_{A^{t}}(C)=\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}>0$
Combined with Theorem 3.2, this contradicts Proposition 3.10, which states that $A^{t}$ is WEF-able.
Case 2: All arcs $i_{j}\to i_{j+1}$ in $C$ have $\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})=\frac{v_{i_{j}}(A_{i_{j}}^{t+1})}{w_{i_{j}}}$. In this case, inequality (1) implies $\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t+1})}{w_{i_{j}}}>\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}$. Notice that all the items in $A_{j_{1}}^{t+1},...,A_{j_{r}}^{t+1}$ are available at iteration $t$, which contradicts the optimality of $\mu^{t}$.
Case 3: Some arcs $i_{j}\to i_{j+1}$ in $C$ satisfy Case 1 and the other arcs satisfy Case 2. Let $l\geq 1$ be the number of arcs in $C$ that satisfy Case 2. We decompose $C$ into a sequence of $l$ edge-disjoint paths, denoted $P_{1},...,P_{l}$, such that the last node of each path is the first node of the next path, and in each path, only the last arc satisfies Case 2. Formally, suppose that some path contains $k\geq 1$ agents, denoted as $i_{1},...,i_{k}$, and $k-1$ arcs. Then for each $21\leq j\leq k-$ 2, $\bar{v}_{i_{j}}\left(A_{i_{j+1}}^{t}\right)=\frac{v_{i_{j}}\left(A_{i_{j+1}}^{t}\right)}{w_{i_{j+1}}}$ and $\bar{v}_{i_{k-1}}\left(A_{i_{k}}^{t}\right)=\frac{v_{i_{k-1}}\left(A_{i_{k-1}}^{t+1}\right)}{w_{i_{k-1}}}$. Since $\overline{cost}_{A^{t}}(C)>0$, there exits a path $P=(i_{1},...,i_{k})$ where $\overline{cost}_{A^{t}}(P)>0$, which implies that:
$0<\sum_{j=1}^{k-1}\left(\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})-\bar{v}_{i_{j}}(A_{i_{j}}^{t})\right)=$
$\sum_{j=1}^{k-2}\left(\frac{v_{i_{j}}(A_{i_{j+1}}^{t})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}\right)+\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t+1})}{w_{i_{k-1}}}-\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t})}{w_{i_{k-1}}}$
The rest of the proof is similar to the proof of Proposition 3.10. We construct another allocation $B^{t}$ randomly as follows:
  1. 1.
    For each agent $11\leq j\leq k-$ 1, we choose one item $o_{i_{j+1}}^{t}$ uniformly from $i_{j+1}$’s bundle and transfer it to $i_{j}$’s bundle.
  2. 2.
    We choose one item $o_{i_{1}}^{t}$ uniformly from $i_{1}$’s bundle to remove.
  3. 3.
    We choose one item $o_{i_{k-1}}^{t+1}$ uniformly from $A_{i_{k-1}}^{t+1}$ and add it to $i_{k-1}$’s bundle.
Thus, the expected change in value between $B^{t}$ and $A^{t}$ is
$\mathbb{E}\left[\sum_{i\in N}\left(v_{i}(B_{i}^{t})-v_{i}(A^{t}_{i})\right)\right]=$
$\sum_{1\leq j\leq k-2}\left(\mathbb{E}\left[v_{i_{j}}(o_{i_{j+1}}^{t})\right]-\mathbb{E}\left[v_{i_{j}}(o_{i_{j}}^{t})\right]\right)+$
$\mathbb{E}\left[v_{i_{k-1}}(o_{i_{k-1}}^{t+1})\right]-\mathbb{E}\left[v_{i_{k-1}}(o_{i_{k-1}}^{t})\right]=$
$\sum_{1\leq j\leq k-2}\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}+\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t+1})}{w_{i_{k-1}}}-\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t})}{w_{i_{k-1}}}$
This is exactly the cost of $P$ which by assumption is greater than 0. However, according to Algorithm 1, $A^{t}$ maximizes the value of an allocation where each agent $i$ receives $w_{i}$ items among the set of $O^{t}$ items. Therefore,
$0\leq\mathbb{E}\left[\sum_{i\in N}\left(v_{i}(B^{t}_{i})-v_{i}(A^{t}_{i})\right)\right]=cost_{A^{t}}(P)$
leading to a contradiction.
To sum up, $A^{t}$ is WEF-able under the original valuations $v$ (with weights $w$), and under the modified valuations $\bar{v}$ (with unit weights). ∎
Proposition 3.15.
For the allocation $A$ computed by Algorithm 1, the subsidy required by an agent given $v$ (with weights $w$) is at most the subsidy required given $\bar{v}$ (with unit weights).
Proof.
Given Observation 3.13, for each $i,j\in N$,
$\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})\geq\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}$
Thus, the cost of any path in the envy graph under the modified function and unit weights is at least the cost of the same path in the weighted envy-graph with the original valuations. ∎
Proposition 3.16.
For the allocation $A$ computed by Algorithm 1, the subsidy to each agent is at most $w_{i}V$ for the modified valuation profile $\bar{v}$.
Proof.
By Proposition 3.14, the allocation $A$ is WEF-able under the valuations $\bar{v}$. Together with Lemma 3.12, if for each $i,j\in N$ it holds that $\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})\geq-V$, the subsidy required for agent $i\in N$ is at most $w_{i}V$ for $\bar{v}$.
$\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})=\sum_{t\in[T]}\bar{v}_{i}(A_{j}^{t})-\sum_{t\in[T]}\bar{v}_{i}(A_{i}^{t})=$
$\sum_{t\in[T-1]}\max\left(\frac{v_{i}(A_{j}^{t})}{w_{j}},\frac{v_{i}(A_{i}^{t+1})}{w_{i}}\right)+\frac{v_{i}(A_{j}^{T})}{w_{j}}-\sum_{t\in[T]}\frac{v_{i}(A_{i}^{t})}{w_{i}}\geq$
$\sum_{t\in[T-1]}\frac{v_{i}(A_{i}^{t+1})}{w_{i}}+\frac{v_{i}(A_{j}^{T})}{w_{j}}-\sum_{t\in[T]}\frac{v_{i}(A_{i}^{t})}{w_{i}}=$
$\frac{v_{i}(A_{j}^{T})}{w_{j}}-\frac{v_{i}(A_{i}^{1})}{w_{i}}\geq-\frac{v_{i}(A_{i}^{1})}{w_{i}}.$
Since $A_{i}^{1}$ contains exactly $w_{i}$ items, $-v_{i}(A_{i}^{1})\geq-w_{i}V$. Hence, $\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})\geq-\frac{w_{i}V}{w_{i}}=-V$. ∎
We are now prepared to prove the main theorem.
Theorem 3.17.
For additive valuations and integer entitlements, Algorithm 1 computes in polynomial time a WEF-able allocation where the subsidy to each agent is at most $w_{i}V$ and the total subsidy is at most $(W-w_{1})V$.
Proof.
Algorithm 1 clearly runs in polynomial-time in $n$ 222it is polynomial in $W$, as the graph $H$ contains $W$ copies. We can compute the max-value matching using reduction to network flow, in time polynomial in $n$..
From Proposition 3.10, we know that $A$ is WEF-able under the original valuations. Combined with Proposition 3.14 and Proposition 3.16 , $A$ is also WEF-able under the modified valuations and requires a subsidy of at most $w_{i}V$ for each agent $i\in N$.
By Proposition 3.15, it follows that under the original valuations, the required subsidy for each agent $i\in N$ is at most $w_{i}V$. From part (4) of Theorem 3.4, there is at least one agent who requires no subsidy, so the required total subsidy is at most $(W-w_{1})V$. ∎
A discussion about the tightness of the bound can be found in appendix A.1.
Brustle et al. [7] proved that the allocation resulting from their algorithm satisfies both EF and EF1. However, in the weighted setup, a WEF-able allocation may not satisfy $WEF(x,y)$ for any $x+y=1$. This also holds for the allocation produced by Algorithm 1 too.
Proposition 3.18.
For general additive valuations, a WEF-able allocation does not necessarily satisfy $WEF(x,y)$ for any $x+y=1$.
Proof.
Suppose 2 agents and 3 items with weights $w_{1}=1,w_{1}=10$ and valuations
$\begin{bmatrix}&1&2&3\\ 1&1&1&1\\ 2&10&10&10\end{bmatrix}$
The allocation $A_{1}=\emptyset,A_{2}=\{o_{1},o_{2},o_{3}\}$ is WEF-able since the cost of the cycle between the two agents in the weighted envy graph is $\frac{3}{10}-0+0-\frac{30}{10}<0$. However, if agent 1 receives at least one item, the cost of the cycle between the two agents in the weighted envy graph becomes at least $\frac{2}{10}-1+10-\frac{20}{10}>0$.
Thus, the only WEF-able allocation is $A_{1}=\emptyset,A_{2}=\{o_{1},o_{2},o_{3}\}$, but this violates WEF$(x,y)$ w.r.t. agent 2. ∎
4 WEF Solutions for Additive Identical Valuations
In this section we assume that all agents have the same additive valuation functions, denoted by $v$.
We present a polynomial-time algorithm for finding a WEF-able allocation with a subsidy bounded by $V$ per agent and a total subsidy bounded by $(n-1)V$. The following example shows that this bound is tight for any weight vector:
Example 4.1.
Consider $n$ agents with integer weights $w_{1}\leq\cdots\leq w_{n}$ and $1+\sum_{i\in N}\left(w_{i}-1\right)$ items all valued at $V$.
To avoid envy, each agent $i$ should receive a total utility of $w_{i}V$, so the sum of all agents’ utilities would be $WV$.
As the sum of all values is $(W-(n-1))V$, a total subsidy of at least $(n-1)V$ is required (to minimize the subsidy per agent, each agent $i\in N$ should receive $w_{i}-1$ items, except for the agent with the highest entitlement (agent $n$), who should receive $W$ items.
The value per unit entitlement of each agent $i<n$ is $V(w_{i}-1)/w_{i}$, and for agent $n$ it is $V$. Therefore, to avoid envy, each agent $i<n$ should receive a subsidy of
$w_{i}\left(1-\frac{w_{i}-1}{w_{i}}\right)V=V$
and the total subsidy required is
$(n-1)V.$
We start by observing that, with identical valuations, the cost of any path in the weighted envy graph is determined only by the agents at the endpoints of that path.
Observation 4.2.
Given an instance with identical valuations, let $A$ be any allocation, and denote by $(i,\ldots,j)$ any path in the envy-graph of $A$ between agents $i,j\in N$. Then,
$cost_{A}(i,\ldots,j)=\sum_{k=i}^{j-1}\left(\frac{v(A_{k+1})}{w_{k+1}}-\frac{v(A_{k})}{w_{k}}\right)=\frac{v(A_{j})}{w_{j}}-\frac{v(A_{i})}{w_{i}}.$
Using Observation 4.2, it is easy to prove the following lemma.
Lemma 4.3.
With identical valuations, every allocation is WEF-able.
Proof.
Consider an allocation $A$ over $m$ items and $n$ agents with identical valuations $v$ and weights $w$. Let $C=(i_{1},...,i_{r})$ be a cycle in $G_{A,w}$. Then,
$\begin{split}&cost_{A}(C)=\\ &\sum_{k=1}^{r-1}\left(\frac{v(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v(A_{i_{1}})}{w_{i_{1}}}-\frac{v(A_{i_{r}})}{w_{i_{r}}}=\\ &\sum_{k=1}^{r}\left(\frac{v(A_{i_{k}}}{w_{i_{k}}}-\frac{v(A_{i_{k}}}{w_{i_{k}}}\right)=0\end{split}$ (2)
Hence, by condition (b) of Theorem 3.2, $A$ is WEF-able. ∎
Our algorithm for finding a WEF-able allocation with bounded subsidy is presented as Algorithm 2.
It first orders the items in $M$ based on their common value, such that $v(o_{1})\geq v(o_{2})\geq...\geq v(o_{m})$.
The algorithm then iteratively traverses the items. At iteration $t$ it selects the agent that minimizes the expression $\frac{v(A_{i}^{t-1}\cup{\{o_{t}\}})}{w_{i}}$, with ties are broken in favor of the agent with the higher index, which corresponds to the agent with the larger $w_{i}$.
Intuitively, this selection minimizes the likelihood that weighted envy is generated. Subsequently, the algorithm allocates item $o_{t}$ to the selected agent.
Example 4.4.
Consider two agents, denoted as $i_{1}$ and $i_{2}$, with corresponding weights $w_{1}=1$ and $w_{2}=\frac{7}{2}$, and three items, namely $o_{1},o_{2},o_{3}$, with valuations $v(o_{1})=v(o_{2})=v(o_{3})=1$, Algorithm 2 is executed as follows:
  1. 1.
    for $t=1$, the algorithm compares $\frac{v(o_{1})}{w_{1}}=1$ and $\frac{v(o_{1})}{w_{2}}=\frac{2}{7}$. Consequently, the algorithm allocates item $o_{1}$ to agent $i_{2}$, resulting in $A_{1}^{1}=\emptyset$ and $A_{2}^{1}=\{o_{1}\}$.
  2. 2.
    for $t=2$, the algorithm compares $\frac{v(o_{2})}{w_{1}}=1$ and $\frac{v(A_{2}^{1}\cup\{o_{2}\})}{w_{2}}=\frac{2}{\frac{7}{2}}=\frac{4}{7}$. Subsequently, the algorithm allocates item $o_{2}$ to agent $i_{2}$, resulting in $A_{1}^{2}=\emptyset$ and $A_{2}^{2}=\{o_{1},o_{2}\}$.
  3. 3.
    for $t=3$, the algorithm compares $\frac{v(o_{3})}{w_{1}}=1$ and $\frac{v(A_{2}^{2}\cup\{o_{3}\})}{w_{2}}=\frac{3}{\frac{7}{2}}=\frac{6}{7}$, Consequently, item $o_{3}$ is allocated to agent $i_{2}$, resulting in $A_{1}^{3}=\emptyset$ and $A_{2}^{3}=\{o_{1},o_{2},o_{3}\}$.
  4. 4.
    agent $i_{1}$ envies agent $i_{2}$ by an amount of $\frac{v(A_{2}^{3})}{w_{2}}-\frac{v(A_{1}^{3})}{w_{1}}=\frac{3}{\frac{7}{2}}=\frac{6}{7}$, and conversely, agent $i_{2}$ envies agent $i_{1}$ by $\frac{v(A_{1}^{3})}{w_{1}}-\frac{v(A_{2}^{3})}{w_{2}}=-\frac{6}{7}$
  5. 5.
    In order to mitigate envy, $s_{i_{1}}=\frac{6}{7}$ and $s_{i_{2}}=0$.
Algorithm 2 Weighted Sequence Protocol For Additive Identical Valuations
1:Instance $(N,M,v,\{w_{i}\}_{i\in N})$ with additive identical valuations.
2:WEF-able allocation $A=(A_{i})_{i\in N}$ with total required subsidy of at most $(n-1)V$.
3:Sort the items based on their value: $v(o_{1})\geq v(o_{2})\geq...\geq v(o_{m})$
4:$A_{i}\leftarrow\emptyset,\forall i\in N$
5:for $t=1$ to $m$ do
6:     $I^{t}\leftarrow\operatorname*{arg\,min}_{i\in N}\left(\frac{v(A_{i}^{t-1}\cup\{o_{t}\})}{w_{i}}\right)$
7:     $i^{t}\leftarrow\max_{i\in I^{t}}\left(i\right)$
8:     $A_{i^{t}}^{t}\leftarrow A_{i^{t}}^{t-1}\cup\{o_{t}\}$
9:     $A_{j}^{t}\leftarrow A_{j}^{t-1}$, for all $j\in N\textbackslash\{i^{t}\}$
10:end for
11:return $A$
Example 4.4 illustrates that the resulting allocation may not meet the WEF(1,0) criterion – Algorithm 2 might allocate all items to the agent with the highest entitlement. However, it does satisfy the WEF(0,1) criterion. The following proposition holds based on the selection rule, which ensures that the cost for any agent $i\in N$ towards the agent who received the new item is bounded by $\frac{V}{w_{i}}$.
Proposition 4.5.
For additive identical valuations, Algorithm 2 computes a $WEF(0,1)$ allocation.
Proof.
We prove by induction that at each iteration $t\in[T]$, where $o_{t}$ is the item assigned in this iteration and $i^{t}$ is the agent receiving this item, $A^{t}$ satisfies $WEF(0,1)$. The claim is straightforward for the first iteration. We assume the claim holds for the $(t-1)$-th iteration and prove it for the $t$-th iteration. Every agent $j\neq i^{t}$ is not envied due to the induction step. For $j=i^{t}$, due to the selection rule, $\frac{v(A^{t}_{i})+v(o^{t})}{w_{i}}=\frac{v(A^{t-1}_{i})+v(o^{t})}{w_{i}}\geq\frac{v(A^{t-1}_{j})+v(o^{t})}{w_{j}}=\frac{v(A^{t}_{j})}{w_{j}}$. This is exactly the definition of WEF$(0,1)$ for the pair $i,j$. ∎
Based on Observation 4.2 and Proposition 4.5, we show that each path starting at agent $i\in N$ under $WEF(0,1)$ allocation is bounded by $\frac{V}{w_{i}}$.
Proposition 4.6.
With identical additive valuations, for every WEF$(0,1)$ allocation $A$, $cost_{A}(\ell_{i}(A))\leq\frac{V}{w_{i}}$, for all $i\in N$.
Proof.
For each agent $i\in N$, denote the most costly path starting at $i$ in that graph by $\ell_{i}(A)=(i,...,j)$ for some agent $j\in N$. Then by Observation 4.2, $cost_{A}(\ell_{i}(A))=\frac{v(A_{j})}{w_{j}}-\frac{v(A_{i})}{w_{i}}$. By definition of WEF$(0,1)$, this latter difference is at most $\frac{v(o)}{w_{i}}$ for some object $o\in A_{j}$. Therefore, the difference is at most $\frac{V}{w_{i}}$. ∎
Theorem 4.7.
For additive identical valuations, Algorithm 2 computes a WEF-able allocation $A$ in $O(mn+m\log(m))$ time, where $\forall i\in N$: $s_{i}\leq V$. Therefore, the total subsidy required is at most $(n-1)V$.
Proof.
Let $A$ be the allocation output by Algorithm 2 under additive identical valuations after $m$ iterations. From Lemma 4.3, it can be deduced that $A$ is WEF-able. Together Proposition 4.5 and Proposition 4.6 imply that, to achieve weighted-envy-freeness under identical additive valuations for the allocation computed by Algorithm 2, the required subsidy per agent $j\in N$ is at most $w_{j}\frac{V}{w_{j}}=V$. In combination with Theorem 3.4, the total required subsidy is at most $(n-1)V$. Note that $W\geq nw_{1}$. Therefore, this bound is better than the one proved in Theorem 3.17: $\left(W-w_{i}\right)V\geq\left(n-1\right)w_{i}V\geq\left(n-1\right)V$.
We now analyze Algorithm 2’s time complexity.
First, sorting the items according the agents’ valuations can be done in $O(m\log(m))$.
Second, the loop in the algorithm runs for $m$ times. where at each iteration $t\in[m]$, finding the set of agents $I^{t}$ and the agent $i^{t}$ within it takes $O(n)$. Also, allocating the item $o_{t}$ to the agent $i^{t}$ takes $O(1)$.
To sum up, Algorithm 2 runs in $O(mn+m\log(m))$. ∎
The upper bound of $(n-1)V$ is tight even for equal entitlements [15]. Interestingly, when either the valuations or the entitlements are identical, the worst-case upper bound depends on $n$, whereas when both valuations and entitlements are different, the bound depends on $W$.
5 WEF Solutions for binary additive Valuations
In this section we focus on the special case of agents with binary additive valuations. We start with a lower bound on the subsidy.
Proposition 5.1.
For every $n$ and every weight vector $\mathbf{w}$, there is an instance with $n$ agents with binary valuations in which the required subsidy in any WEF allocation is at least $\frac{W}{w_{2}}-1$.
Proof.
There is one item. Agents $1$ and $2$ value the item at $1$ and the others at 00. If agent $i\in\{1,2\}$ gets the item, then the other agent $j$ must get subsidy $\frac{w_{j}}{w_{i}}$. To ensure that other agents do not envy $j$’s subsidy, every other agent $k\not\in\{1,2\}$ must get subsidy $\frac{w_{k}}{w_{i}}$. The total subsidy is $\frac{W}{w_{i}}-1$. The subsidy is minimized by giving the item to agent $2$, since $w_{2}\geq w_{1}$. This gives a lower bound of $\frac{W}{w_{2}}-1$. ∎
We now show how to compute a WEF-able allocation where the subsidy given to each agent $i\in N$ is at most $\frac{w_{i}}{w_{1}}V=\frac{w_{i}}{w_{1}}$.
In the case of binary valuations, Algorithm 1 is inefficient in three ways: (1) the maximum-cost matching does not always prioritize agents with higher entitlements, (2) there may be situations where an agent prefers items already allocated in previous iterations, while the agent holding those items could instead take unallocated ones, and (3) the algorithm works only for agents with integer weights. To illustrate these issues, see the following example:
Example 5.2.
Consider two agents with corresponding weights $w_{1}=1$ and $w_{2}=2$, and five items. The valuation functions are outlined below:
$\begin{bmatrix}&o_{1}&o_{2}&o_{3}&o_{4}&o_{5}\\ i_{1}&1&1&1&1&1\\ i_{2}&1&1&1&1&0\end{bmatrix}$
One possible scenario is that Algorithm 1 proceeds as follows:
  1. 1.
    At the first iteration, agent 1 receives $o_{1}$ and agent 2 receives $o_{2},o_{3}$.
  2. 2.
    At the second iteration, agent 1 receives $o_{4}$, as this matching has a total value of $1$, which is the highest possible value at this stage;
  3. 3.
    Similarly,at the third iteration, agent 1 receives $o_{5}$.
In this case, the subsidy required by agent 2 is $\left(\frac{2}{1}-\frac{2}{2}\right)2=2$ which is higher than $W-w_{2}=3-2=1$. An allocation requiring less subsidy would be $A_{1}=\{o_{5}\}$ and $A_{2}=\{o_{1},o_{2},o_{3},o_{4}\}$. Here, the subsidy required by agent 1 is $(\frac{4}{2}-1)2=1$. (an even better allocation would be $A_{1}=\{o_{4},o_{5}\}$ and $A_{2}=\{o_{1},o_{2},o_{3}\}$, which is already WEF without any subsidy).
We address them by modifying the General Yankee Swap (GYS) algorithm introduced by Viswanathan et al. in [26].
GYS starts with an initial allocation, denoted $A^{0}$. In this allocation all agents’ bundles are empty, $A_{i}^{0}=\emptyset$, $\forall i\in N$. For convenience, we add a dummy agent $i_{0}$ and assume that all items are initially assigned to $i_{0}$: $A_{i_{0}}^{0}=M$. The algorithm runs in $T$ iterations . For $t\in\{0,\ldots,T\}$, we denote by $A^{t}$ the allocation at the end of the iteration $t$. Throughout this algorithm, we divide the agents into two sets:
  1. 1.
    $U^{t}$: The agents remaining in the game at the beginning of the iteration $t$.
  2. 2.
    $N\setminus U^{t}$: The agents who were removed from the game in earlier iteration $t^{\prime}<t$. Agents are removed from the game when the algorithm deduces that their utility cannot be improved.
As long as not all the objects have been allocated, at every iteration $t\in[T]$, the algorithm looks for the agents maximizing the gain function (Viswanathan et al. in [26]) among $U^{t}$, i.e., the agents remaining in the game at this iteration.
In our setting we use the gain function: $\left(\frac{w_{i}}{v_{i}(A_{i}^{t})+1},1\right)$, which means that we first select the agents with the minimal potential to increase envy towards them (the maximum $\frac{w_{i}}{v_{i}(A_{i}^{t})+1}$), and then break ties in favor of the agent with the higher index, which corresponds to the agent with the larger $w_{i}$.
The selected agent then faces a choice: either acquire an unallocated item or take an item allocated to another agent. Regardless of the choice, the agent’s utility increases by 1. If the agent decides to take an allocated item from another agent, the affected agent must then decide whether to acquire an unallocated item or take an allocated item from a third agent to maintain their utility, and so on. This results in a path from agent $i$ to the dummy agent $i_{0}$, where items are passed between agents until an unallocated item is reached.
We call this a transfer paths. When an agent is selected, the algorithm attempts to find a transfer path from this agent to $i_{0}$, maintaining utilities for all agents except the initiator, whose utility increases by 1. If no such path exists, the agent will no longer receive any items, and the agent is removed from the game.
We use the polynomial-time method by Viswanathan et al. in [26] to find transfer paths.
Algorithm 3 differs from $GYS$ in the following way: at the beginning of iteration $t$, the algorithm first removes all agents without a transfer path originating from them (Phase 1). Then, it selects an agent based on the gain function to allocate a new item to that agent (Phase 2).
Algorithm 3 Weighted Sequence Protocol For Additive Binary Valuations
1:Instance $(N,M,\{v_{i}\}_{i\in N},\{w_{i}\}_{i\in N})$ with additive binary valuations.
2:WEF-able allocation $A=(A_{i})_{i\in N}$ with total required subsidy of at most $\frac{W}{w_{1}}-1$.
3:$A_{i_{0}}^{0}\leftarrow M$, and $A_{i}^{0}\leftarrow\emptyset$ for each $i\in N$triangleright All items initially are unassigned
4:$U^{1}\leftarrow N$
5:$t\leftarrow 1$
6:while $U^{t}\neq\emptyset$ do
7:
8:Phase 1 – Removing all agents without a transfer path starting from them
9:
10:     for $i\in U^{t}$ do
11:         Find a transfer path starting at $i$
12:         if  no such path exists then
13:              $U^{t}\leftarrow U^{t}\setminus\{i\}$
14:         end if
15:     end for
16:
17:Phase 2 – Updating the resulting allocation
18:
19:     $I^{t}\leftarrow\arg\max_{i\in U^{t}}\left(\frac{w_{i}}{v_{i}(A_{i}^{t-1})+1}\right)$triangleright Choose the agent who maximizes the gain function
20:     $i^{t}\leftarrow\max_{i\in S}\left(i\right)$triangleright Break ties using index
21:     Find the shortest transfer path starting at $i^{t}$
22$\triangleright$ this path must exist, as $i^{t}$ would have been removed otherwise
23:     Transfer the items along the path and update the allocation $A^{t}$
24:     $U^{t+1}\leftarrow U^{t}$
25:     $t\leftarrow t+1$
26:end while
27:return $A^{t}$
Example 5.3.
For Example 5.2, the algorithm is executed as follows:
  1. 1.
    For $t=1$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$, $\frac{1}{w_{2}}=\frac{1}{2}$. Consequently, the algorithm searches for the shortest transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{1}$ to agent $i_{2}$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{1}=\emptyset$ and $A_{2}^{1}=\{o_{1}\}$.
  2. 2.
    For $t=2$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$ and $\frac{v(A_{2}^{1})+1}{w_{2}}=\frac{2}{2}$. Since those values are equal, the algorithm compares $w_{1}=1$ and $w_{2}=2$. Subsequently, the algorithm searches for the shortest transfer path starting at $i_{1}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{2}$ to agent $i_{2}$, yielding $A_{1}^{2}=\emptyset$ and $A_{2}^{2}=\{o_{1},o_{2}\}$.
  3. 3.
    For $t=3$, the algorithm compares $\frac{1}{w_{1}}=1$ and $\frac{v_{2}(A_{2}^{2})+1}{w_{2}}=\frac{3}{2}$. As a result, the algorithm searches for the shortest transfer path starting at $i_{1}$ and ending at $i_{\text{0}}$, and finds the path $(i_{1},i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $i_{1}$, producing $A_{1}^{3}=\{o_{3}\}$ and $A_{2}^{3}=\{o_{1},o_{2}\}$.
  4. 4.
    For $t=4$, the algorithm compares $\frac{v_{1}(A_{1}^{3})+1}{w_{1}}=2$ and $\frac{v_{2}(A_{2}^{3})+1}{w_{2}}=\frac{3}{2}$. Thus, the algorithm searches for the shortest transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{4}$ to agent $i_{2}$, leading $A_{1}^{4}=\{o_{3}\}$ and $A_{2}^{4}=\{o_{1},o_{2},o_{4}\}$.
  5. 5.
    For $t=5$, the algorithm compares $\frac{v_{1}(A_{1}^{4})}{w_{1}}=2$ and $\frac{v_{2}(A_{2}^{2})+1}{w_{2}}=2$. Since those values are equal, the algorithm compares $w_{1}=1$ and $w_{2}=2$. As a consequence, the algorithm searches for the shortest transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{1},i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $i_{2}$ from $i_{1}$’s bundle and the item $o_{5}$ to agent $i_{1}$ from $i_{\text{0}}$’s bundle. The final allocation is $A_{1}^{5}=\{o_{5}\}$ and $A_{2}^{5}=\{o_{1},o_{2},o_{3},o_{4}\}$.
  6. 6.
    The envies are stated below:
    $\begin{bmatrix}&i_{1}&i_{2}\\ i_{1}&0&\frac{4}{2}-\frac{1}{1}=1\\ i_{2}&0-\frac{4}{2}=-2&0\\ \end{bmatrix}$
  7. 7.
    In order to mitigate envy, $s_{i_{1}}=1$ and $s_{i_{2}}=0$. Note that the agent with second-smallest weight, $i_{2}$, receives a subsidy of 00.
Definition 5.4.
(Viswanathan et al. [26]) An allocation $A$ is said to be non-redundant if for all $i\in N$, we have $v_{i}(o)=|A_{i}|$.
That is, $v_{j}(A_{i})\leq|A_{i}|=v_{i}(A_{i})$ for every $i,j\in N$. Similarly to Lemma 4.3, we can prove the following Lemma.
Lemma 5.5.
With binary valuations, every non-redundant allocation is WEF-able.
Proof.
Consider a non-redundant allocation $A$ over $m$ items and $n$ agents with binary additive valuations $v$ and weights $w$. Let $C=(i_{1},...,i_{r})$ cycle in $G_{A,w}$. Then,
$\begin{split}&cost_{A}(C)=\\ &\sum_{k=1}^{r-1}\left(\frac{v_{i_{k}}(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v_{i_{r}}(A_{i_{1}})}{w_{i_{1}}}-\frac{v_{i_{r}}(A_{i_{r}})}{w_{i_{r}}}\leq^{\text{due to non-redundancy}}\\ &\sum_{k=1}^{r-1}\left(\frac{v_{i_{k+1}}(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v_{i_{1}}(A_{i_{1}})}{w_{i_{1}}}-\frac{v_{i_{r}}(A_{i_{r}})}{w_{i_{r}}}=\\ &\sum_{k=1}^{r}\left(\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)=0\end{split}$ (3)
Hence, by condition (b) of Theorem 3.2, $A$ is WEF-able. ∎
The following is an analogue of Observation 4.2:
Observation 5.6.
Let $A$ be any non-redundant allocation. Let $L=(i,\ldots,j)$ be a path in $G_{A,w}$. Then $cost_{A}(L)\leq\frac{|A_{j}|}{w_{j}}-\frac{|A_{i}|}{w_{i}}$.
Lemma 3.1 in [26] shows that the allocation produced by GYS is non-redundant. The same is true for our variant:
Lemma 5.7.
In any iteration $t\in[T]$ of Algorithm 3, the allocation $A^{t}$ is non-redundant.
Proof.
We prove by induction that at each iteration $t\in[T]$, $A^{t}$ remains non-redundant.
For the base case, $A^{0}$ is an empty allocation and is therefore non-redundant. Now, assume that at iteration $t-1$, $A^{t-1}$ is non-redundant. Let $A^{t}$ be the allocation resulting from iteration $t$.
If $A^{t-1}=A^{t}$, meaning no agent received a new item, the process is complete. Otherwise, let $i^{t}$ the agent who receives new item. Agent $i^{t}$ obtains an item via the transfer path $L=(i^{t}=i_{1},\ldots,i_{k})$. For each $1\leq j<k$, agent $i_{j}$ receives the item $o_{j}$ from the bundle of $i_{j+1}$, given that $v_{i_{j}}(o_{j})=1$. Agent $i_{k}$ receives a new item $o_{k}$ from the bundle of $i_{0}$, with $v_{i_{k}}(o_{k})=1$.
Additionally, for each $1<j\leq k$, item $o_{j-1}$ is removed from agent $i_{j}$’s bundle where $v_{i_{j}}(o_{j-1})=1$, since $A^{t-1}$ is non-redundant.
For agents not on the transfer path $L$, their bundles remain unchanged. Thus, for each agent $i\in N$, it holds that $v_{i}(A_{i}^{t})=|A_{i}^{t}|$, confirming that $A^{t}$ is non-redundant. ∎
Our main task is to prove the following:
Theorem 5.8.
For additive binary valuations, Algorithm 3 computes a WEF-able allocation where the subsidy to each agent $i\in N$ is at most $\frac{w_{i}}{w_{1}}$ in polynomial-time. Moreover, the total subsidy is bounded by $\frac{W}{w_{1}}-1$.
Based on Lemma 5.7 and Lemma 5.5 it is established that for every iteration $t\in[T]$, $A^{t}$ is both non-redundant and WEF-able. Therefore, the remaining task is to establish subsidy bounds.
We focus on two groups: $U^{t}$ and $N\setminus{U^{t}}$. The selection rule simplifies setting limits for the first group (Proposition 5.10). However, understanding the dynamics of the second group presents challenges, as the selection rule is not applicable for them. Hence, the primary objective is to determine the bounds for these subsidies (Proposition 5.14).
Lemma 5.9.
Let $j\in N$ be any agent. After Phase 1 at iteration $t$, if $i\in U^{t}$, then $\frac{|A^{t}_{j}|}{w_{j}}-\frac{|A^{t}_{i}|}{w_{i}}\leq\frac{1}{w_{i}}.$
5.9 is proved similarly to the proof of 4.5, based on the selection rule.
Proof.
If $j$ has never been selected to receive an item, then $|A^{t}_{j}|=0$ and the lemma is trivial.
Otherwise, let $t^{\prime}\leq t$ be the latest iteration in which $j$ was selected. By the selection rule, $\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}$. Then by non-redundancy, $\frac{|A_{j}^{t^{\prime}}|}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}=\frac{v_{i}(A_{i}^{t^{\prime}})}{w_{i}}+\frac{1}{w_{i}}=\frac{|A_{i}^{t^{\prime}}|}{w_{i}}+\frac{1}{w_{i}}$.
As $|A_{j}^{t}|=|A_{j}^{t^{\prime}}|$ and $|A_{i}^{t}|\geq|A_{i}^{t^{\prime}}|$, the lemma follows. ∎
From Lemma 5.9, we can conclude the following:
Proposition 5.10.
After Phase 1 at iteration $t$, if $i\in U^{t}$, then $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{i}}$.
Proof.
Assume $\ell_{i}(A^{t})=(i=i_{1},...,i_{k})$ is the path with the highest total cost starting at $i$ in the weighted envy graph corresponding to the allocation at the end of the iteration, $G_{A^{t},w}$.
Observation 5.6 implies $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{|A_{i_{k}}^{t}|}{w_{i_{k}}}-\frac{|A_{i}^{t}|}{w_{i}}$.
As $i\in U^{t}$, Lemma 5.9 implies $\frac{|A_{i_{k}}^{t}|}{w_{i_{k}}}-\frac{|A_{i}^{t}|}{w_{i}}\leq\frac{1}{w_{i}}$. ∎
Proposition 5.10 ensures an upper bound on the subsidy required for agents remaining in the game: for each agent $i\in U^{t}$ and iteration $t\in T$, $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{i}}\leq\frac{1}{w_{1}}$.
Note that, in the special case of identical binary valuations, all agents remain in the game simultaneously. Therefore, 5.10 ensures an upper bound of $w_{i}\cdot\frac{1}{w_{i}}=1$ on the subsidy of each agent, which implies a total subsidy of at most $n-1$. However, this subsidy bound is not applicable for agents who were removed from the game. For an agent $i$ removed from the game, we can prove a subsidy bound of $w_{i}\cdot\frac{1}{w_{j}}$, where $j$ is some other agent that remains in the game.
To prove this upper bound, we need to establish several claims about agents who are removed from the game. First, we show that an agent removed from the game does not desire any item held by an agent who remains in the game (5.11). As a result, these removed agents will not be included in any transfer path (5.12).
Next, we demonstrate that if the cost of a path originating from these removed agents at some iteration $t$ exceeds the cost at $t^{\prime}\leq t$ — the iteration when the agent was removed — then there exists an arc in this path, $(i_{j},i_{j+1})$, such that $v_{i}(A_{j}^{t})=0$ (5.13). Based on these claims, we prove that if at iteration $t$, the cost of the maximum-cost path starting from agent removed from the game at $t^{\prime}<t$ exceeds its cost at $t^{\prime}$, we can upper-bound it by $\frac{1}{w_{1}}$.
Proposition 5.11.
Let $i$ be an agent removed from the game in Phase 1 at iteration $t^{\prime}$. Then for all $j\in U^{t^{\prime}}$ who were not removed in Phase 1 at iteration $t^{\prime}$, $v_{i}(A^{t^{\prime}}_{j})=0$.
Moreover, for all $t>t^{\prime}$ and all $j\in U^{t}$, $v_{i}(A^{t}_{j})=0$.
Proof.
Suppose that, at iteration $t^{\prime}$, after Stage 2, $v_{i}(A_{j}^{t^{\prime}})\neq 0$. This implies there exists some item $o\in A_{j}^{t^{\prime}}$ such that $v_{i}(o)=1$. We consider two cases:
  1. 1.
    $o\in A_{j}^{t^{\prime}-1}$. In this case, at the beginning of iteration $t^{\prime}$, there exists a transfer path from $i$ to $j$. Moreover, there is a transfer path from $j$ to $i_{0}$ at iteration $t^{\prime}$ (otherwise, $j$ would have been removed from the game at $t^{\prime}$ as well). Concatenating these paths gives a transfer path from $i$ to $i_{0}$.
  2. 2.
    $o\not\in A_{j}^{t^{\prime}-1}$, that is, $j$ received item $o$ during iteration $t^{\prime}$, from some other agent $j^{\prime}$ (where $j^{\prime}=i_{0}$ is possible). At the beginning of iteration $t^{\prime}$, there exists a transfer path from $i$ to $j^{\prime}$. Moreover, there is a transfer path from $j^{\prime}$ to $i_{0}$, which is used to transfer the newly allocated item. Concatenating these paths gives a transfer path from $i$ to $i_{0}$.
Both cases contradict the assumption that $i$ removed at $t^{\prime}$.
To prove the claim for $t>t^{\prime}$, we use induction over $t$.
We assume the claim holds for iteration $t-1>t^{\prime}$ and prove it for iteration $t$. Assume, contrary to the claim, that there exists an agent $i$ who was removed at iteration $t^{\prime}$, and an agent $j\in U^{t}$ such that $v_{i}(A_{j}^{t})\neq 0$. By the induction hypothesis and the fact that $U^{t-1}\supseteq U^{t}$, we have $v_{i}(A_{j}^{t-1})=0$. Therefore, during iteration $t$, $j$ must have received a new item $o_{j}$ that $i$ values at $1$.
If $o_{j}$ was part of $i_{0}$’s bundle at $t^{\prime}$, then the transfer path starting at $i$, $(i,i_{0})$, must have already existed at $t^{\prime}$.
Alternatively, if $o_{j}$ was originally in another agent’s bundle at $t^{\prime}$, say agent $k\in N$, then there must have been an iteration between $t^{\prime}$ and $t$ in which $o_{j}$ has been transferred from $k$ to another agent, while agent $k$ is compensated by some other item $o_{k}$, that agent $k$ wants.
If $o_{k}$ was part of $i_{0}$’s bundle at $t^{\prime}$, then the transfer path starting at $i$, $(i,k,i_{0})$, must have existed at $t^{\prime}$. Otherwise, $o_{k}$ was in another agent’s bundle at $t^{\prime}$, and it, too, would be transferred to a different bundle in a later iteration.
Since the number of items is finite, this process must eventually lead to an item that was in $A_{i_{0}}^{t^{\prime}}$, forming a transfer path starting at $i$ at iteration $t^{\prime}$ — a contradiction.∎
Proposition 5.12.
Let $i$ be an agent removed from the game in Stage 1 at iteration $t^{\prime}$. Then, for all $t\geq t^{\prime}$, $i$ will not be included in any transfer path.
Proof.
First, note that $v_{i}(A_{i_{0}}^{t})=0$; otherwise, $i$ would not have been removed at $t^{\prime}$. Next, any agent $j$, who receives an item from $i_{0}$’s bundle in iteration $t$, must be in $U^{t}$. By Proposition 5.11, $v_{i}(A_{j}^{t})=0$, which means that $i$ can not receive any item from $j\in U^{t}$ in compensation for another item from their own bundle. Thus, any transfer path includes only agents in $U^{t}$. ∎
Proposition 5.13.
Consider an iteration $t$ and an agent $i\notin U^{t}$, who was removed from the game in Phase 1 at iteration $t^{\prime}<t$. Let $L=(i=i_{1},\ldots,i_{k})$ be a path in the weighted envy graph starting at $i$. If $cost_{A^{t}}(L)>cost_{A^{t^{\prime}-1}}(L)$, then there must exist $j\in\{1,\ldots,k-1\}$ such that $i_{j}\notin U^{t}$, $i_{j+1}\in U^{t}$ and $v_{i_{j}}(A_{i_{j+1}}^{t})=0$.
Proof.
Let $t>t^{\prime}$ be the earliest iteration in which $cost_{A^{t}}(L)>cost_{A^{t^{\prime}-1}}(L)$. This implies that there is at least one agent, say $11\leq j^{\prime}\leq k-$ 1, such that agent $i_{j^{\prime}+1}$ has received a new item that $i_{j^{\prime}}$ desires. In other words, $i_{j^{\prime}+1}$ was part of a transfer path at iteration $t$, and by Proposition 5.12, $i_{j^{\prime}+1}\in U^{t}$ (in particular, $j^{\prime}\geq 2$).
Since $i_{1}\notin U^{t}$ and $i_{j^{\prime}+1}\in U^{t}$, there must exist some $11\leq p<j^{\prime}+$ 1 such that $i_{p}\notin U^{t}$ and $i_{p+1}\in U^{t}$. By Proposition 5.11, $v_{i_{p}}(A^{t}_{i_{p+1}})=0$. ∎
Proposition 5.14.
Let $i\notin U^{t}$, be an agent who was removed from the game at iteration $t^{\prime}<t$. Then, at the end of iteration $t$, $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{1}}$.
Proof.
Recall that $\ell_{i}(A^{t^{\prime}-1})$ is the highest-cost path starting from $i$ at iteration $t^{\prime}-1$ (before agent $i$ removed). In particular, $cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t}))\leq cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t^{\prime}-1}))$. Moreover, by Observation 5.6 we have $cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t^{\prime}-1}))\leq\frac{|A_{i_{k}}^{t^{\prime}-1}|}{w_{i_{k}}}-\frac{|A_{i}^{t^{\prime}-1}|}{w_{i}}$ when $i_{k}$ is the last agent in $\ell_{i}(A^{t^{\prime}-1})$. Combined with Lemma 5.9, this gives $cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t^{\prime}-1}))\leq\frac{1}{w_{i}}$. Therefore, if $cost_{A^{t}}(\ell_{i}(A^{t}))\leq cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t}))$, then we are done.
Assume now that $cost_{A^{t}}(\ell_{i}(A^{t}))>cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t}))$. Denote the path $\ell_{i}(A^{t})$ by $(i=i_{1},\ldots,i_{k})$.
From Proposition 5.13, there exists $j\in\{1,\ldots,k-1\}$ such that $i_{j}\notin U^{t}$, $i_{j+1}\in U^{t}$ and $v_{i_{j}}(A_{i_{j+1}}^{t})=0$. Then, by Observation 5.6:
$cost_{A^{t}}(\ell_{i}(A^{t}))=$ (4)
$cost_{A^{t}}(i,...,i_{j})+cost_{A^{t}}(i_{j},i_{j+1})+cost_{A^{t}}(i_{j+1},...,i_{k})\leq$
$\leq\left(\frac{|A_{i_{j}}^{t}|}{w_{i_{j}}}-\frac{|A_{i}^{t}|}{w_{i}}\right)+\left(0-\frac{|A_{i_{j}}^{t}|}{w_{i_{j}}}\right)+cost_{A^{t}}(\ell_{i_{j+1}}(A^{t}))\leq$
$cost_{A^{t}}(\ell_{i_{j+1}}(A^{t}))$ (5)
Since $i_{j+1}\in U^{t}$, it follows from (5) and Proposition 5.10 that
$cost_{A^{t}}(\ell_{i}(A^{t}))\leq cost_{A^{t}}(\ell_{i_{j+1}}(A^{t}))\leq\frac{1}{w_{i_{j+1}}}\leq\frac{1}{w_{1}}.$
We are now prepared to establish the proof of Theorem 5.8.
Proof.
Together Proposition 5.10 and Proposition 5.14 establish that for every $i\in N$ and $t\in[T]$, $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{1}}$. Along with Lemma 5.7 and Lemma 5.5, Algorithm 3 computes a WEF-able allocation $A^{T}$ where the required subsidy per agent $i\in N$ is at most $w_{i}\cdot\frac{1}{w_{1}}=\frac{w_{i}}{w_{1}}$.
As there is at least one agent who requires no subsidy (see Theorem 3.4), the total required subsidy is at most $\frac{W-w_{1}}{w_{1}}=\frac{W}{w_{1}}-1$.
We complete the proof of Theorem 5.8 by demonstrating that Algorithm 3 runs in polynomial-time. We represent the valuations using a binary matrix $X$ where $v_{i}(o_{j})=1\Longleftrightarrow X(i,j)=1$. Hence, the allocation of items to the bundle of $i_{0}$ at line 1 can be accomplished in $O(mn)$ time.
At each iteration $t$ of the while loop, either $A_{i_{0}}$ or $U^{t}$ reduced by 1, ensuring that the loop runs at most $m+n$ times.
Let $T_{v}$ represent the complexity of computing the value of a bundle of items, and $T_{\phi}$ denote the complexity of computing the gain function. Both are polynomial in $m$.
According to Viswanathan et al. [26], finding a transfer path starting from agent $i\in N$ (or determining that no such path exists) takes $O(T_{v}\log m)$. Removing agents at the start of each iteration incurs a complexity of $O(nT_{v}\log m)$. Furthermore, as stated in Viswanathan et al. [26], identifying $i^{t}$ requires $O(nT_{v})$. Updating the allocation based on the transfer path, according to the same source, takes $O(m)$.
Thus, each iteration has a total complexity of $O(nT_{v}\log m+nT_{v}+T_{v}\log m+m)=O\left(nT_{v}\log m+m\right)$.
In conclusion, Algorithm 3 runs in $O\left(\left(m+n\right)\left(nT_{v}\log m+m\right)\right)$, which is polynomial in both $m$ and $n$. ∎
A discussion about the tightness of the bound can be found in appendix A.2
Notice that since the output allocation from Algorithm 3 is non-redundant, $A$ maximizes the USW . Moreover, Example 5.3, is not WEF(1,0) (No matter which item is removed from $i_{2}$’s bundle, $i_{1}$ will still feel envious). Hence we can conclude that a $MUSW^{M,N,v}$ allocation is not necessarily WEF(1,0) either. However, it is WEF(0,1).
Proposition 5.15.
For additive binary valuations, Algorithm 3 computes a $WEF(0,1)$ allocation.
Proof.
We prove by induction that at each iteration $t\in[T]$ $A^{t}$ satisfies $WEF(0,1)$. This means that for every $i,j\in N$, there exists a set of items $B\subseteq A_{j}^{t}$ of size at most $1$ such that $\frac{v_{i}(A_{i}^{t})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j}^{t})}{w_{j}}$.
The claim is straightforward for the first iteration. We assume the claim holds for the $(t-1)$-th iteration and prove it for the $t$-th iteration. Note that $v_{i}(A_{i}^{t-1})\leq v_{i}(A_{i}^{t})$.
  1. 1.
    $A_{j}^{t}=A_{j}^{t-1}$ and $A_{i}^{t}=A_{i}^{t-1}$: the claim holds due to the induction step.
  2. 2.
    $v_{i}(A_{j}^{t})=v_{i}(A_{j}^{t-1})$: This is the case where $j$ was not included in a transfer path, or was included but was not the first agent in the path, and exchanged an item for a new one, both having the same value for $i$. By the induction assumption, there exists some singleton $B^{t-1}\subseteq A_{j}^{t-1}$ such that $\frac{v_{i}(A_{i}^{t-1})+v_{i}(B^{t-1})}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}$. There exists some singleton $B^{t}\subseteq A_{j}^{t}$, with $v_{i}(B^{t})=v_{i}(B^{t-1})$. Hence, $\frac{v_{i}(A_{i}^{t})+v_{i}(B^{t})}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+v_{i}(B^{t-1})}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}$.
  3. 3.
    $v_{i}(A_{j}^{t})<v_{i}(A_{j}^{t-1})$: This is the case where $j$ was included in a transfer path but exchanged an item $i$ values for an item $i$ does not value. Then $\frac{v_{i}(A_{i}^{t})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}>\frac{v_{i}(A_{j}^{t})}{w_{j}}$ for a set $B\subseteq A_{j}^{t}$ of size at most $1$.
  4. 4.
    $v_{i}(A_{j}^{t})>v_{i}(A_{j}^{t-1})$: there are two subcases:
    1. (a)
      If $j$ is the first agent in the transfer path and received a new item $o$ such that $v_{i}(o)=1$, then $\frac{v_{i}(A_{i}^{t-1})+1}{w_{i}}\geq\frac{v_{j}(A_{j}^{t-1})+1}{w_{j}}$ due to the selection rule, and $\frac{v_{j}(A_{j}^{t-1})+1}{w_{j}}\geq\frac{v_{i}(A_{j}^{t-1})+1}{w_{j}}$ due to non-redundancy.
      We can conclude that $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+1}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})+1}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}$. The claim holds for $B=\{o\}\subseteq A_{j}^{t}$.
    2. (b)
      If $j$ was not the first agent in the path, but exchanged an item that $i$ does not value for an item $o$ that $i$ values, $v_{i}(o)=1$. Let $t^{\prime}<t$ represent the most recent iteration in which agent $j$ was selected and received a new item. Note that $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}\geq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}\geq\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}$ due to the selection rule.
      Assume to the contrary that $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}<\frac{v_{i}(A_{j}^{t})}{w_{j}}$. Then,
      $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}<\frac{v_{i}(A_{j}^{t})}{w_{j}}\leq\frac{v_{j}(A_{j}^{t})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq$
      $\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}\leq\frac{v_{i}(A_{i}^{t})+1}{w_{i}},$
      a contradiction. Hence, $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}\geq\frac{v_{i}(A_{j}^{t})}{w_{j}}$ and the claim holds for $B=\{o\}\subseteq A_{j}^{t}$.
6 Conclusions And Future Work
We studied the problem of subsidy minimization required for achieving weighted envy-freeness among agents with varying entitlements when allocated indivisible items.
Previous work in the unweighted setting of subsidies relied on basic characterizations of EF that fail in the weighted settings. This poses interesting challenges in the new setting.
We show that an allocation is WEF-able only if its weighted envy graph does not contain positive cost cycles. Unlike the unweighted scenario ([15]), a given allocation that maximizes the utilitarian welfare across all reassignments of its bundles to agents is not a condition for WEF-able. The gap between the weighted setting and the unweighted setting in EF characterization raises the open question of other conditions under which weighted EF allocations may exist.
We’ve shown polynomial-time algorithms to compute WEF-able allocations for general, identical, and binary additive valuations in the weighted setting. The proved subsidy bounds are $\left(W-w_{1}\right)V$, $(n-1)V$, and $\frac{W}{w_{1}}-1$, respectively.
While our $(n-1)V$ bound is tight, some gap remains for $\left(W-w_{1}\right)V$ and $\frac{W}{w_{1}}-1$ for the weighted setting. However for identical and binary additive valuations, our bounds align with those of the unweighted setting. Like the previous literature, we focus on additive valuations. This highlights the need for further investigation into refining subsidy bounds for the general additive case and extending results to non-additive valuations in both weighted and unweighted contexts.
APPENDIX
Appendix A Tightness of the Subsidy Bounds
A.1 Subsidy Bound of Algorithm 1
As Theorem 3.17 implies, Algorithm 1 computes a WEF-able allocation with a total subsidy of at most $(W-w_{1})V$. However, this bound is not tight. To understand why, consider the case of $2$ items, each valued at $V$ by agent $i\in\{1,\ldots,n-1\}$, who has an entitlement of $w_{i}\geq 2$, and $V-\epsilon$ by all other agents. Our algorithm will allocate all the items to agent $i$, resulting in a subsidy of $\frac{w_{j}}{w_{i}}2(V-\epsilon)$ by each other agent $j\neq i\in N$, leading to a total subsidy of $(W-w_{i})\frac{2(V-\epsilon)}{w_{i}}$, for arbitrarily small $\epsilon>0$.
In general, a WEF-able allocation can achieve a lower subsidy by allocating one item to another agent with higher index $j>i$, i.e, $w_{j}\geq w_{i}$. For instance, if one item is allocated to such agent $j$, agent $i$ envies agent $j$ by $\frac{V}{w_{j}}-\frac{V}{w_{i}}\leq 0$, and agent $j$ envies agent $i$ by $\frac{V-\epsilon}{w_{i}}-\frac{V-\epsilon}{w_{j}}<\frac{2(V-\epsilon)}{w_{i}}$. If $\frac{V-\epsilon}{w_{i}}-\frac{V-\epsilon}{w_{j}}\leq 0$, then no subsidy is required. Otherwise, the subsidy required by agent $j$ is $\left(\frac{V-\epsilon}{w_{i}}-\frac{V-\epsilon}{w_{j}}\right)w_{j}<\frac{w_{j}}{w_{i}}\cdot 2(V-\epsilon)$. The subsidy required by each other agent $k\neq i,j$ is significantly lower than $w_{k}\cdot\frac{w_{j}}{w_{j}}\cdot\frac{2(V-\epsilon)}{w_{i}}=\frac{w_{k}}{w_{i}}\cdot 2(V-\epsilon)$. Therefore, the required total subsidy is significantly lower than $(W-w_{i})\frac{2(V-\epsilon)}{w_{i}}$.
In both cases, the resulting total subsidy bound is better than the bound obtained by allocating all items to agent $i$.
A.2 Subsidy Bound of Algorithm 3
As Theorem 5.8 implies, Algorithm 3 computes a WEF-able allocation with a total subsidy of at most $\frac{W}{w_{1}}-1$. However, this bound is not necessary tight.
First, there are cases where breaking ties in favor of the agent with the minimum entitlement leads to a WEF-able allocation with a better subsidy bound.
Example A.1.
Consider three agents with weights $w_{1}=1,w_{2}=2,w_{3}=3$ and three items. Agent 1 and agent 3 value all the items at 1, while agent 2 values all the items at 0.
the algorithm is executed as follows:
  1. 1.
    For $t=1$, since agent $2$ values all items at 0, the algorithm removes this agent from the game. Then, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1},\frac{1}{w_{3}}=\frac{1}{3}$. Consequently, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{1}$ to agent $3$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{1}=\emptyset,A_{2}^{1}=\emptyset$ and $A_{3}^{1}=\{o_{1}\}$.
  2. 2.
    For $t=2$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$ and $\frac{v_{3}(A_{3}^{1})+1}{w_{3}}=\frac{2}{3}$. Subsequently, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{2}$ to agent $3$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{2}=\emptyset,A_{2}^{2}=\emptyset$ and $A_{3}^{2}=\{o_{1},o_{2}\}$.
  3. 3.
    For $t=3$, the algorithm compares $\frac{1}{w_{1}}=1$ and $\frac{v_{3}(A_{3}^{2})+1}{w_{3}}=\frac{3}{3}=1$. Since those values are equal, the algorithm compares $w_{1}=1$ and $w_{3}=3$. As a result, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $3$, producing $A_{1}^{3}=\emptyset,A_{2}^{3}=\emptyset$ and $A_{3}^{3}=\{o_{1},o_{2},o_{3}\}$.
In the resulting allocation, agent 1 envies agent 3 by $1$, resulting in a subsidy of $1$ by agent 1. Agent 2 now envies agent 1 by 1 due to the subsidy, resulting in a subsidy of $2$ by agent 2. The total subsidy required is $3$. However, if at $t=3$ we allocate an item to agent 1, resulting in the allocation $A_{1}=\{o_{1}\},A_{2}=\emptyset,A_{3}=\{o_{2},o_{3}\}$, agent 3 envies agent 1 by $1-\frac{2}{3}=\frac{1}{3}$, which means the subsidy required by agent 3 is $1$. Agent 2 now envies agent 3 by $\frac{1}{3}$, resulting in a subsidy of $\frac{2}{3}$ by agent 2. The total subsidy required is $\frac{5}{3}<3$.
This is not yet clear in which cases it is preferable to break ties in favor of higher or lower entitlements.
Secondly, if agent 1, who has an entitlement of $w_{1}$, is the only agent who requires 0 subsidy, then the subsidy required by each other agent $j\neq 1\in N$ is not necessary bounded by $\frac{w_{j}}{w_{1}}$.
Example A.2.
Consider two agent with weight $w_{1}=2,w_{2}=3$ and two items, both values at 1 by both agents. the algorithm is executed as follows:
  1. 1.
    For $t=1$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1},\frac{1}{w_{3}}=\frac{1}{3}$. Consequently, the algorithm searches for the shortest transfer path starting at $3$ and ending at $i_{\text{0}}$, and finds the path $(3,i_{\text{0}})$. The algorithm transfers the item $o_{1}$ to agent $3$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{1}=\emptyset$ and $A_{2}^{1}=\{o_{1}\}$.
  2. 2.
    For $t=2$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{2}$ and $\frac{v_{2}(A_{2}^{1})+1}{w_{3}}=\frac{2}{3}$. Subsequently, the algorithm searches for the shortest transfer path starting at $2$ and ending at $i_{\text{0}}$, and finds the path $(2,i_{\text{0}})$. The algorithm transfers the item $o_{2}$ to agent $2$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{2}=\{o_{2}\}$ and $A_{2}^{2}=\{o_{1}\}$.
In the resulting allocation, agent 2 envies agent 1 by $\frac{1}{2}-\frac{1}{3}=\frac{1}{6}$, resulting in a subsidy of $\frac{1}{2}$ by agent 3. That is, agent 1 with lower entitlement is the only agent required 0 subsidy, but the subsidy required by agent 2 is smaller than $\frac{w_{2}}{w_{1}}=\frac{3}{2}$.
We still do not know how to tighten this bound further for the general case.